### v1

# Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models

Abstract
We develop an empirical Bayes (EB) G-modeling framework for short-panel linear models with multidimensional heterogeneity and nonparametric prior. Specifically, we allow heterogeneous intercepts, slopes, dynamics, and a non-spherical error covariance structure. We establish identification and consistency of the nonparametric maximum likelihood estimator (NPMLE) under general conditions, and provide low-level sufficient conditions for several models of empirical interest. Conditions for regret consistency of the resulting EB estimators are also established. The NPMLE is computed using a Wasserstein-Fisher-Rao gradient flow algorithm adapted to panel regressions. Using data from the Panel Study of Income Dynamics, we find that the slope coefficient for potential experience is substantially heterogeneous and negatively correlated with the random intercept, and that error variances and autoregressive coefficients vary significantly across individuals. The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates.
Keywords: G-modeling, nonparametric maximum likelihood, shrinkage estimation, Wasserstein-Fisher-Rao gradient flow, income dynamics
1 Introduction
Understanding differences in individual behavior is the primary goal of many economic analyses. Why do workers with similar experience exhibit different earnings profiles? Why do economic agents with similar observed characteristics respond differently to economic shocks and policy interventions? A challenge for empirical researchers in addressing these questions is the presence of individual-level differences that affect behavior but are not captured by observed variables, commonly referred to as unobserved heterogeneity. While it is not hard to allow for unit-specific intercepts in linear regressions, entertaining heterogeneous slope parameters and innovation variances is more challenging, and in a frequentist setting, this usually requires parametric assumptions to gain tractability at the cost of robustness.
This paper considers a setting in which we observe $\{(Y_{it},X_{2,it}):t=1,\ldots,T\}$ for each unit $i=1,\ldots,N$, assuming they are independent and identically distributed (i.i.d.) across $i$. We propose a framework for short-$T$ linear panel models with heterogeneous intercepts, slope coefficients, and error distributions. To fix ideas, consider the dynamic linear panel model with an exogenous covariate $X_{2,it}$ and AR(1) errors:
$Y_{it}$ $=a_{i}+b_{i}X_{2,it}+u_{it},$ (1a)
$u_{it}$ $=\rho_{i}u_{it-1}+\sigma_{i}e_{it},\quad e_{it}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1),\qquad t\geq 1,$ (1b)
where $a_{i}\in\mathbb{R}$ is a random intercept, $b_{i}\in\mathbb{R}$ a random slope, $\rho_{i}\in(-1,1)$ a random autoregressive coefficient, $\sigma_{i}^{2}>0$ a random variance, and $u_{i0}$ follows the stationary distribution $\mathcal{N}(0,\sigma_{i}^{2}/(1-\rho_{i}^{2}))$ conditional on $(\sigma_{i}^{2},\rho_{i})$. We refer to the model in (1) as the HIVDX model since it admits heterogeneity in four dimensions: intercepts ($I$), error variances ($V$), dynamics ($D$) through the autoregressive coefficient, and the coefficient on the covariate $X_{2}$. The widely studied Gaussian location-scale model $Y_{it}=a_{i}+\sigma_{i}e_{it}$, as well as models that impose $\rho_{i}=\rho$ and $b_{i}=b$, are special cases of HIVDX.
We analyze the HIVDX model and the special cases in Section 2 using a general framework that partitions the random parameters into two sets: regression coefficients $\beta_{i}=(a_{i},b_{i})$ of dimension $d_{\beta}=2$, and covariance parameters $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$ of dimension $d_{\delta}=2$. The parameters can be estimated separately for each $i$ by maximum likelihood. Instead, we assume that the random parameters $\theta_{i}\equiv(\beta_{i},\delta_{i})$ are drawn from a true but unknown common distribution $G_{*}$. Viewing $G_{*}$ as a prior, the optimal decision rule for each unit under compound squared-error loss is the posterior mean ${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\mid Y_{i},X_{2,i}]$, where $Y_{i}=(Y_{i1},\ldots,Y_{iT})^{\prime}$ is a $T\times 1$ vector of outcomes, $X_{2,i}=(X_{2,i1},\ldots,X_{2,iT})^{\prime}$ is a vector of observed covariates, and $\mathbb{E}_{G_{*}}[\cdot]$ denotes expectation with respect to $\theta_{i}\sim G_{*}$. We refer to $\theta_{i}^{*}$ as the oracle estimator.
Specifying that $\theta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}G_{*}$ is grounded in the principle of empirical Bayes (EB) introduced by Robbins (1956). This approach is intrinsically frequentist because it estimates the prior distribution from the data rather than specifying it. EB estimation improves on individual-level maximum likelihood estimation through shrinkage, borrowing strength across units to learn both the individual parameters $\theta_{i}$ and their population distribution $G_{*}$. Following the terminology of Efron (2014) the procedure is referred to as G-modeling when $G_{*}$ is estimated directly. The resulting EB estimator is given by $\hat{\theta}^{\mathrm{EB}}_{i}:=\mathbb{E}_{\hat{G}}[\theta_{i}\,|\,Y_{i},X_{2,i}]$, where $\hat{G}$ denotes an estimator of $G_{*}$. An EB estimator $\hat{\tau}_{i}^{{\mathrm{EB}}}(\theta_{i})=\mathbb{E}_{\hat{G}}[\tau_{i}\,|\,Y_{i},X_{2,i}]$ for general unit-specific parameters $\tau_{i}$ defined as functions of $\theta_{i}$ can also be defined. While parametric G-modeling is convenient, nonparametric modeling guards against misspecification. Recent reviews of EB and its applications in economics include Koenker and Gu (2024), Koenker and Gu (2025), and Walters (2024). To our knowledge, no nonparametric EB analysis has considered all four sources of heterogeneity present in the HIVDX model.
We study identification, estimation and and optimality of the HIVDX model within an EB framework. The generality of our model with $d_{\beta}>1$ and $d_{\delta}>0$ introduces several challenges. First, while assuming the identification of $G_{*}$ in the Gaussian location model with i.i.d. errors without formal proof may be defensible, this assumption cannot be taken for granted in the more complex settings. Theorem 1 establishes generally that identification requires: (i) the design matrix to have full column rank within its support, and (ii) the covariance structure to remain identifiable after removing the mean component by appropriately differencing the data. These conditions imply that $d_{\beta}<T$, leaving $T-d_{\beta}$ degrees of freedom to identify the covariance structure, thereby imposing a constraint on the number of heterogeneous variance parameters ($d_{\delta}$). Proposition 1 applies Theorem 1 to the location model ($d_{\beta}=1$) and illustrates how differencing facilitates identification. Specifically, we show that $T\geq 3$ suffices for AR(1) errors, whereas ARMA(1,1) errors require $T\geq 4$. The analysis is, however, more involved when $d_{\beta}>1$, as the mean and variance components are convolved within the outcome distribution. Now whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariate. Proposition 3 establishes that if $X_{2,it}$ is such that $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0$ or $\Delta X_{2,i2}\neq\Delta X_{2,i4}$, where $\Delta$ denotes the first-difference operator, then $G_{*}$ is nonparametrically identified with $T=4$. However, Proposition 4 shows that identification is possible with $T\geq 5$ for a different $X_{2,it}$ satisfying the weaker condition that $\Delta X_{2,it}\neq 0$ for some $t=2,\ldots,T$.
The second challenge lies in establishing the consistency of the HIVDX model. The classic proof of Kiefer and Wolfowitz (1956) assumes compactness, but when $\beta_{i}$ is allowed to be unbounded, the space of mixing distributions is non-compact. While the case of ${d_{\beta}}=1$ can be resolved by embedding $\mathbb{R}$ into its compact completion $\infty$$\infty$], the multidimensional setting (${d_{\beta}}\geq 2$) presents unique difficulties because mass can diverge along infinitely many distinct directions. To address this, Section 4 adopts a metric induced by the vague topology to effectively compactify the space of prior distributions. Theorem 2 establishes the almost sure consistency of the NPMLE with respect to this metric. Building on this result, Theorem 3 in Section 5 provides conditions for the regret consistency (equivalently, asymptotic optimality) of the EB estimator for $\tau_{i}$, taking into account that $\tau_{i}$ may involve functions of the random coefficients and the data. For example, the oracle one-step-ahead prediction under quadratic loss is ${\tau}^{*}_{i}=\mathbb{E}_{G_{*}}[a_{i}+b_{i}X_{2,iT+1}+\rho_{i}Y_{iT}\,|\,Y_{i},X_{i},X_{2,iT+1}]-\mathbb{E}_{G_{*}}[\rho_{i}a_{i}+\rho_{i}b_{i}X_{2,iT}\,|\,Y_{i},X_{i}]$. The prediction error $\hat{\tau}^{{\mathrm{EB}}}_{i}-\tau_{i}^{*}$ thus involves interaction of the errors in estimating the random coefficients and $(Y_{i},X_{i})$. Proposition 5 provides low-level sufficient conditions for the EB prediction to be regret-consistent.
The third challenge arises from the computational complexity associated with multiple sources of heterogeneity. Building on the interior-point method of Koenker and Mizera (2014), Gu and Koenker (2017) discretize $G_{*}$ over a fixed grid and solve the dual formulation of the NPMLE via convex optimization. While effective in one or two dimensions, these methods do not scale well as the dimension of the parameter space increases. We also approximate $G_{*}$ using a discrete distribution, but we employ an algorithm capable of flexibly handling higher-dimensional heterogeneity. To this end, Section 6 adapts the Wasserstein-Fisher-Rao (WFR) gradient flow algorithm of Yan, Wang, and Rigollet (2024) to the panel regression setting. In this algorithm, the Wasserstein step adaptively relocates support points to accelerate convergence, while the Fisher-Rao step updates the probability weights via Bayes’ rule.
We use the EB approach to analyze the income data in the Panel Study of Income Dynamics (PSID) in Section 7. The model allows earnings $Y_{it}$ to have unit-specific mean $a_{i}$ and response $b_{i}$ to experience $X_{2,it}$, with a heteroskedastic error covariance structure that depends on internal dynamics $\rho_{i}$ and exposure to shocks through $\sigma_{i}^{2}$. We find that (i) the slope coefficient $b_{i}$ for potential experience exhibits substantial heterogeneity and is negatively correlated with the random intercept $a_{i}$, and (ii) there is pronounced cross-sectional heterogeneity in both the variance ($\sigma_{i}^{2}$) and the dynamics ($\rho_{i}$). The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates. Section 8 presents results from Monte Carlo experiments calibrated to the application.
Before turning to the main part of the paper, we briefly discuss recent work that, while not directly related to our approach, are conceptually connected to the broader EB literature. Chen (2024) analyzes conditionally Gaussian settings with known variances $\sigma_{i}^{2}$, constructing studentized data by estimating the conditional mean and variance of outcomes given $\sigma_{i}^{2}$, and then estimating the prior for use in EB decision rules. The paper emphasizes the importance of nonparametric estimation in both stages and illustrates the method using data on economic mobility. In contrast, we allow for unobserved $\sigma_{i}^{2}$ and heterogeneous dynamics across multiple time periods.
Kwon (2025) considers a multivariate model in which $Y_{i}$ given $\theta_{i}\in\mathbb{R}^{T}$ follows $\mathcal{N}(\theta_{i},\Sigma_{i})$ with known covariance matrices $\Sigma_{i}$. Under the hierarchical specification $\theta_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\Lambda)$, the shrinkage estimator $\Lambda(\Lambda+\Sigma_{i})^{-1}Y_{i}$ is shown to be optimal when $\Lambda$ minimizes Stein’s unbiased risk estimate, without requiring normality for validity. Our framework differs in that $\Sigma_{i}$ is unobserved and we allow for random slopes on regressors, whereas Kwon (2025) assumes homogeneous coefficients and removes the regression component prior to estimation. Cheng, Ho, and Schorfheide (2025) propose a parametric EB estimator for two-way effects in linked data, using a prior that captures patterns of assortative matching. Their framework is more related to Kwon (2025) than ours in that the shrinkage parameters are chosen by minimizing an unbiased risk estimate. Both studies are applied to teacher value-added models.
Gaillac (2024) develops posterior mean estimators for cross-sectional random coefficient models in which the coefficients are independent of the regressors. In a panel extension, the paper constructs time averages for each unit and allows heterogeneity only in time-invariant regressors, while assuming homogeneous coefficients for time-varying regressors. In contrast, our framework focuses on random coefficients associated with regressors that vary over time.
Adusumilli, Gu, and Tao (2025) investigate EB methods in compound adaptive experiments, where the outcome of each arm follows a normal distribution with an unknown mean. They demonstrate that the risk guarantees for G-modeling, originally derived under i.i.d. sampling, continue to hold for adaptively collected data without knowledge of the sampling algorithm. In contrast, F-modeling is shown to yield biased estimates in this setting.
Recent work has extended nonparametric EB methods beyond the classical Gaussian location model to a variety of settings, including heteroskedastic models (e.g., Soloff, Guntuboyina, and Sen, 2025), high-dimensional or multivariate regression (e.g., Mukherjee, Sen, and Sen, 2023; Fan, Guan, Shen, and Wu, 2023; Jiang and Guntuboyina, 2025), and variance estimation (e.g., Ignatiadis and Sen, 2025). These works only consider a static model while our focus is the heterogeneous dynamic panel model presented above as HIVDX.
2 The Econometric Setup and EB Modeling
We consider a panel of observations on $N$ cross-sectional units over $T$ time periods, focusing on a short-panel framework in which $N$ grows while $T$ remains fixed. Our main setup is a general heterogeneous coefficient (HC) model that nests the HIVDX model:
$Y_{i}=X_{i}\beta_{i}+P_{i}e_{i},$ (2)
where $Y_{i}=(Y_{i1},\ldots,Y_{iT})^{\prime}$ is a $T\times 1$ vector of outcomes, $X_{i}=(X_{i1},\ldots,X_{iT})^{\prime}$ is a $T\times d_{\beta}$ matrix of observed covariates, and $e_{i}=(e_{i1},\ldots,e_{iT})^{\prime}$ is a vector of independent Gaussian errors. We assume that the first column of $X_{i}$ is $\iota_{T}$, the $T$-dimensional vector of ones. For each $i$, $P_{i}$ is lower triangular, and the corresponding conditional covariance matrix $\Sigma_{i}\equiv P_{i}P_{i}^{\prime}$ is positive definite. To capture its dependence on heterogeneous variance parameters $\delta_{i}\in\mathbb{R}^{{d_{\delta}}}$, we write $P_{i}=P(\delta_{i})$, where $\delta\mapsto P(\delta)$ is a continuously differentiable function of $\delta$.
To see that the HIVDX model is a special case of the HC model for particular $(\beta_{i},\delta_{i})$, note that since $e_{it}$ are i.i.d. $\mathcal{N}(0,1)$, $u_{i}=(u_{i1},\ldots,u_{iT})^{\prime}$ is normally distributed with conditional covariance matrix $\Sigma_{i}$, whose $(t,s)$ entry is given by
$\Sigma_{i,ts}=\sigma_{i}^{2}\frac{\rho_{i}^{|t-s|}}{1-\rho_{i}^{2}},\quad t,s=1,\ldots,T.$ (3)
Hence, $P_{i}$ is defined via the Cholesky decomposition of $\Sigma_{i}=P_{i}P_{i}^{\prime}$ with $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$. We develop our theoretical framework through the lens of the HC model since it provides a unified way to analyze a broader class of dynamics such as ARMA errors. Let $I_{T}$ denote the $T\times T$ identity matrix. The following assumptions will be used throughout.
Assumption 1 (DGP).
Let $(Y_{i},X_{i})_{i=1}^{N}$ be identically and independently distributed (i.i.d.) data, satisfying (2). In addition, the following conditions hold.
  1. (i)
    The pairs $(\beta_{i},\delta_{i})_{i=1}^{N}$ are i.i.d. draws from an unknown distribution $G_{*}$ and are independent of $(X_{i})_{i=1}^{N}$.
  2. (ii)
    The parameter space for $\delta_{i}$ is a compact subset $\mathcal{K}_{\delta}\subseteq\mathbb{R}^{{d_{\delta}}}$. For each $\delta_{i}\in\mathcal{K}_{\delta}$, the covariance matrix $\Sigma_{i}\equiv P(\delta_{i})P(\delta_{i})^{\prime}$ satisfies $\underaccent{\bar}{c}\,I_{T}\leq\Sigma_{i}\leq\bar{c}\,I_{T}$ for some $0<\underaccent{\bar}{c}<\bar{c}<\infty$.
  3. (iii)
    $e_{i}\sim\mathcal{N}(0,I_{T})$.
Conditional on $(\beta_{i},\delta_{i})$ and $X_{i}$, the outcome vector $Y_{i}$ follows the distribution $\mathcal{N}(X_{i}\beta_{i},\Sigma_{i})$. Individuals in this model are distinguished by types characterized by their heterogeneous features. Given the information structure, individual types cannot be directly uncovered from the data, but their distributions can be identified.
Let $\theta_{i}:=(\beta_{i},\delta_{i})$ collect all heterogeneous parameters, lying in the space $\Theta:=\mathbb{R}^{{d_{\beta}}}\times\mathcal{K}_{\delta}\subseteq\mathbb{R}^{{d_{\theta}}}$ with ${d_{\theta}}={d_{\beta}}+{d_{\delta}}$. If $\theta_{i}$ were observed, the conditional density of $Y_{i}$ given $X_{i}$ is
$\ell(Y_{i}\,|\,X_{i},\theta_{i}):=\frac{1}{({2\pi})^{T/2}|P(\delta_{i})|}\exp\left(-\frac{1}{2}\|P(\delta_{i})^{-1}(Y_{i}-X_{i}\beta_{i})\|^{2}\right).$
When $\theta_{i}$ is unobserved but its distribution is given as $G$, the marginal density of $Y_{i}$ given $X_{i}$, obtained by integrating out $\theta_{i}$ from the likelihood, is
$f_{G}(Y_{i},X_{i}):=\int_{\Theta}\ell(Y_{i}\,|\,X_{i},\theta)dG(\theta),$
where the independence between $\theta_{i}$ and $X_{i}$ is used. [^1] We use $G_{*}$ to denote the unknown true distribution and $G$ to denote a generic distribution in the class of possible distributions that includes $G_{*}$. The notation $f_{G}(Y_{i},X_{i})$ emphasizes that the marginal density of $Y_{i}$ given $X_{i}$ depends on $G$.
For the estimation of $(\theta_{i})_{i=1}^{N}$, two conceptual frameworks are available. The separate decision framework treats each $\theta_{i}$ as fixed and minimizes the individual risk $\mathbb{E}[\|\hat{\theta}_{i}-\theta_{i}\|^{2}\,|\,\theta_{i}]$. This yields the individual MLE,
$\hat{\theta}^{\operatorname{MLE}}_{i}:=\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\theta\in\Theta}\ell(Y_{i}\,|\,X_{i},\theta),$
which provides an asymptotically efficient estimator for $\theta_{i}$ as $T\to\infty$. However, Stein (1956) and James and Stein (1961) demonstrate that the MLE is inadmissible under squared-error loss when the dimension of the parameter vector is at least three (here, $N\geq 3$). In this setting, the MLE is dominated by a shrinkage estimator that achieves lower total risk by pooling information across observations.
The second approach, the compound decision framework, treats $(\theta_{i})_{i=1}^{N}$ as i.i.d. draws from a common distribution $G_{*}$ and seeks to minimize the compound risk $\mathbb{E}_{\theta_{i}\sim G_{*}}[\|\hat{\theta}_{i}-\theta_{i}\|^{2}]$. Viewing $G_{*}$ as a prior, the optimal decision rule for each unit is the posterior mean, defined in the introduction as the oracle estimator:
${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\,|\,Y_{i},X_{i}]=\frac{\int_{\Theta}\theta\ell(Y_{i}\,|\,X_{i},\theta)dG_{*}(\theta)}{f_{G_{*}}(Y_{i},X_{i})}.$ (4)
This estimator improves efficiency by pooling information across units, thereby reducing the variance relative to the individual maximum likelihood estimators $\hat{\theta}^{\operatorname{MLE}}_{i}$. These efficiency gains come at the cost of introducing individual bias (i.e., $\mathbb{E}[{\theta}^{*}_{i}\,|\,\theta_{i}]\neq\theta_{i}$), but the reduction in the overall compound mean squared error dominates this cost. In essence, ${\theta}^{*}_{i}$ shrinks $\hat{\theta}^{\operatorname{MLE}}_{i}$ toward the prior.
In practice, the compound decision rule is infeasible due to the unknown $G_{*}$. In Bayesian analysis, the prior $G_{*}$ is typically specified as a parametric distribution, either by fixing its hyperparameters or by assigning them their own distributions. Efron and Morris (1973) show that the James-Stein shrinkage estimator can be interpreted as a parametric EB estimator, with $G_{*}$ serving as a parametric prior whose hyperparameters are estimated from the data.
An Empirical Bayes analysis can approximate the infeasible oracle decision rule in one of two ways: F-modeling or G-modeling. F-modeling directly computes the shrinkage term from the data without explicit knowledge of $G_{*}$. For example, assuming $\Sigma_{i}\equiv\Sigma$ for some known homogeneous covariance matrix $\Sigma$ (that is, $\rho_{i}=\rho$ and $\sigma_{i}^{2}=\sigma^{2}$ for some known $\rho$ and $\sigma^{2}$), the (infeasible) F-modeling EB estimator is
$\hat{\beta}^{\operatorname{F-EB}}_{i}:=\hat{\beta}^{\operatorname{MLE}}_{i}+(X_{i}^{\prime}\Sigma^{-1}X_{i})^{-1}\frac{\partial}{\partial\hat{\beta}}\log\hat{f}_{\hat{\beta}|X}(\hat{\beta}^{\operatorname{MLE}}_{i},X_{i}),$ (5)
where $\hat{\beta}^{\operatorname{MLE}}_{i}=(X_{i}^{\prime}\Sigma^{-1}X_{i})^{-1}X_{i}^{\prime}\Sigma^{-1}Y_{i}$ and $\hat{f}_{\hat{\beta}|X}(\hat{\beta},x)$ is an estimator for $f_{\hat{\beta}|X}(\hat{\beta},x)$ – the conditional density of $\hat{\beta}^{\operatorname{MLE}}_{i}$ given $X_{i}=x$ under $G_{*}$. In practice, $\Sigma$ is typically unknown and must also be estimated. Equation (5), commonly known as Tweedie’s formula in the literature, shows that the posterior mean of $\beta_{i}$ adjusts the corresponding MLE by a shrinkage term derived from the cross-sectional distribution of the MLE. [^2] In the simplest case when $Y\,|\,\theta\overset{\text{ind.}}{\sim}\mathcal{N}(\theta,1)$ with $T=1$, the Tweedie formula is $\mathbb{E}_{G_{*}}[\theta|y]=y+\tfrac{d}{dy}\log f_{G_{*}}(y)$, where $f_{G_{*}}(\cdot)$ denotes the marginal density of $Y$ under $G_{*}$. The magnitude of the shrinkage term grows with the variance of $\hat{\beta}^{\operatorname{MLE}}_{i}$, as measured by $(X_{i}^{\prime}\Sigma^{-1}X_{i})^{-1}$, and with the informativeness of the prior, as reflected in $f_{\hat{\beta}|X}$. [^3] It can be shown that the oracle-decision errors and the shrinkage term are uncorrelated, that is, $\operatorname{Cov}_{G_{*}}({\beta}^{*}_{i}-\beta_{i},\hat{\beta}^{\operatorname{MLE}}_{i}-{\beta}^{*}_{i})=0$, where ${\beta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\beta_{i}\,|\,Y_{i},X_{i}]$, implying that the variance reduction is attributable to the shrinkage term in (5).
F-modeling typically relies on the assumption that the likelihood belongs to the linear exponential family, offering computational convenience when this condition holds. Walters (2024) provides a survey of F-modeling applications within labor economics. Liu, Moon, and Schorfheide (2020) estimate the posterior mean of $a_{i}$ by F-modeling using Tweedie’s formula. However, this method does not readily extend to the HIVDX model. We adopt the alternative of G-modeling, which explicitly estimates the unknown prior $G_{*}$. Shen and Wu (2025) demonstrate that G-modeling has superior statistical properties in Poisson mixture models than F-modeling.
While parametric G-modeling imposes a specific functional form on the prior, nonparametric G-modeling characterizes the true distribution $G_{*}$ as the maximizer of the population log-likelihood:
$G_{*}\in\operatornamewithlimits{arg\hskip 1.00006ptmax}_{G\in\mathcal{G}}F(G),$ (6)
where $\mathcal{G}$ is the class of all probability distributions on $\Theta$ and $F(G):=\mathbb{E}[\log f_{G}(Y_{i},X_{i})]$ denotes the expected marginal log-likelihood (i.e., the negative risk). This formulation motivates estimating $G_{*}$ via its sample analog:
$\hat{G}\in\operatornamewithlimits{arg\hskip 1.00006ptmax}_{G\in\mathcal{{G}}}F_{N}(G),$ (7)
where the empirical criterion is given by $F_{N}(G):=N^{-1}\sum_{i=1}^{N}\log f_{G}(Y_{i},X_{i})$. We refer to $\hat{G}$ as the nonparametric maximum likelihood estimator (NPMLE) of $G_{*}$. Substituting $\hat{G}$ for $G_{*}$ in (4) yields the G-modeling EB estimator for $\theta_{i}$:
$\hat{\theta}_{i}^{{\mathrm{EB}}}:=\frac{\int_{\Theta}\theta\ell(Y_{i}\,|\,X_{i},\theta)d\hat{G}(\theta)}{\int_{\Theta}\ell(Y_{i}\,|\,X_{i},\theta)d\hat{G}(\theta)}.$ (8)
The NPMLE is built on the same principle as parametric MLE, namely that $G_{*}$ is the maximizer of the population log-likelihood defined in (6). If $G_{*}$ is the unique maximizer of $F$ over $\mathcal{G}$, it is said to be (point-)identified. However, it is known that the NPMLE in (7) may not be unique in multidimensional settings; in such cases, $\hat{G}$ refers to any maximizer. Thus, before presenting an algorithm to compute $\hat{G}$ in Section 6, we first establish conditions for the identification of $G_{*}$ in Section 3, and then show in Section 4 that consistent estimation is possible despite multiple sources of heterogeneity.
3 Identification
This section establishes identification of the HC model defined in (2). Let $\mathcal{G}$ denote the class of all distributions on $\Theta\equiv\mathbb{R}^{{d_{\beta}}}\times\mathcal{K}_{\delta}$, defined as
$\mathcal{G}:=\mathcal{P}(\mathbb{R}^{{d_{\beta}}}\times\mathcal{K}_{\delta})=\{G\in\mathcal{P}(\mathbb{R}^{d_{\theta}}):\operatorname{supp}_{G}(\delta_{i})\subseteq\mathcal{K}_{\delta}\},$ (9)
where $\mathcal{K}_{\delta}\subseteq\mathbb{R}^{{d_{\delta}}}$ is a compact subset in which $\delta_{i}$ is supported, as in Assumption 1. Since any solution to the MLE in (6) must lie in the set
$\left\{G\in\mathcal{G}:f_{G}(Y_{i},X_{i})=f_{G_{*}}(Y_{i},X_{i})\ \ \text{a.s.}\right\},$ (10)
the identified set of distributions consists of distributions consistent with the observed marginal likelihood. We now state the conditions on the covariates and the parametric form of $P(\cdot)$ required for the identification of $G_{*}$.
Assumption 2.
There exists a pair $({\boldsymbol{X}},\boldsymbol{M})$ of fixed matrices with ${\boldsymbol{X}}\in\operatorname{supp}(X_{i})$ and $\boldsymbol{M}\in\mathbb{R}^{T\times(T-{d_{\beta}})}$ such that the following conditions hold:
  1. (i)
    $\operatorname{rank}({\boldsymbol{X}})={d_{\beta}}\leq T-1$, $\operatorname{rank}(\boldsymbol{M})=T-{d_{\beta}}\geq 1$, and $\boldsymbol{M}^{\prime}{\boldsymbol{X}}=0$.
  2. (ii)
    For any $\delta,\tilde{\delta}\in\mathcal{K}_{\delta}$,
    $\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}=\boldsymbol{M}^{\prime}P(\tilde{\delta})P(\tilde{\delta})^{\prime}\boldsymbol{M}\quad\text{if and only if}\quad\delta=\tilde{\delta}$.
Assumption 2(i) requires $\boldsymbol{X}$} to have full column rank, paralleling the familiar rank condition in regression. Identifying $\beta_{i}$ is generally infeasible unless some ${\boldsymbol{X}}\in\operatorname{supp}(X_{i})$ has full column rank. This ensures the existence of an annihilator matrix $\boldsymbol{M}$ with rank $T-{d_{\beta}}$. [^4] The choice of $\boldsymbol{M}$ is unique (up to a nonsingular linear transformation) given $\boldsymbol{X}$}, since a different choice of $\boldsymbol{M}$ can be written as $\tilde{\boldsymbol{M}}=\boldsymbol{M}A$ for some nonsingular $A$. We further require that ${d_{\beta}}\leq T-1$ so that $\boldsymbol{M}$ has non-null rank and can be used in part (ii). This restriction is needed because, in our setting, the covariances $\Sigma_{i}$ depend on unknown $\delta_{i}$.
Assumption 2(ii) imposes that $\delta\mapsto\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$ is a one-to-one function of $\delta\in\mathcal{K}_{\delta}$, meaning that the variance parameter remains identifiable after removing the effect of ${\boldsymbol{X}}\beta_{i}$. The intuition is as follows. Since the concatenated matrix $({\boldsymbol{X}},\boldsymbol{M})$ is of full rank, the information in $Y_{i}$, conditional on $X_{i}={\boldsymbol{X}}$, can be expressed as
$\left[\begin{matrix}\boldsymbol{M}^{\prime}Y_{i}\\ {\boldsymbol{X}}^{\prime}Y_{i}\end{matrix}\right]=\left[\begin{matrix}0\\ {\boldsymbol{X}}^{\prime}{\boldsymbol{X}}\beta_{i}\end{matrix}\right]+\left[\begin{matrix}\boldsymbol{M}^{\prime}P_{i}e_{i}\\ {\boldsymbol{X}}^{\prime}P_{i}e_{i}\end{matrix}\right],$
where $\boldsymbol{M}^{\prime}Y_{i}$ is a Gaussian mixture with mean zero and variance $\boldsymbol{M}^{\prime}P_{i}P_{i}^{\prime}\boldsymbol{M}=\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$. The distribution of $\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$ can therefore be identified from the distribution of $\boldsymbol{M}^{\prime}Y_{i}$. Assumption 2(ii) states that $\delta_{i}$ is identified since $\delta\mapsto\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$ is one-to-one.
To identify $G_{*}$, the remaining step is to recover the conditional distribution of $\beta_{i}$ given $\delta_{i}$. Let $\mathcal{E}\in\mathbb{R}^{T\times(T-{d_{\beta}})}$. The key idea is to express
$\boldsymbol{M}^{\prime}Y_{i}+\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ $=(\boldsymbol{M}^{\prime}+\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime})P(\delta_{i})e_{i}+\mathcal{E}^{\prime}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})\beta_{i}$
as a perturbation of $\boldsymbol{M}^{\prime}Y_{i}$ with noise $\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ injected. We measure the shift in the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}+\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ relative to the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}$ identified earlier. This comparison reveals the conditional distribution of $\mathcal{E}^{\prime}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})\beta_{i}$ given $\delta_{i}$ as $\mathcal{E}$ traverses $\mathbb{R}^{T\times(T-{d_{\beta}})}$, which then identifies $\beta_{i}$ given $\delta_{i}$ via the Cramer-Wold device.
The following theorem formalizes this result. The proof of Theorem 1 as well as the proofs of all other theorems and propositions are given in Appendix A.
Theorem 1
Under Assumptions 1 and 2, $G_{*}$ is identified in the class $\mathcal{G}=\mathcal{P}(\mathbb{R}^{{d_{\beta}}}\times\mathcal{K}_{\delta})$.
The primary challenge in identifying $G_{*}$ is to distinguish the mean and variance components that are convolved into the outcome distribution. For instance, assume that $Y_{i}\sim\mathcal{N}(0,1)$ is a scalar observation. The observed distribution of $Y_{i}$ can be rationalized by a continuum of models: $Y_{i}=a_{i}+\sigma_{i}e_{i}$, where $a_{i}\sim\mathcal{N}(0,v_{a})$ and $\sigma_{i}\equiv\sqrt{1-v_{a}}$ for $v_{a}\in[0,1]$. Thus, one cannot tell which model generates the observed distribution without imposing further restrictions on $a_{i}$ or $\sigma_{i}$. Bruni and Koch (1985) establish general identification results for Gaussian mixtures under a compact support assumption, which requires both the mean and variance components to be compactly supported. [^5] These results have recently attracted attention in the identification of latent distributions in learning models within labor economics (see, e.g., Pastorino (2024); Bunting, Diegert, and Maurel (2024); de Paula, Gualdani, Pastorino, and Salgado (2025)). This branch of the literature does not overlap with the broader econometrics literature on the identification of mixture models (see Compiani and Kitamura (2016) for a review). Using a type of identification-at-infinity argument, they show that both components can be separated by leveraging variation in the tails of $Y_{i}$. However, this argument breaks down as soon as the mean component places a slight nonzero mass in the tails, producing observationally indistinguishable yet fundamentally different models, as illustrated above. In contrast, our identification strategy rests on the rank and one-to-one restrictions that preclude this type of identification failure.
Theorem 1 extends to models that include homogeneous slope parameters:
$Y_{i}=X_{i}\beta_{i}+W_{i}\gamma_{*}+P_{i}e_{i},$ (11)
where $W_{i}=(W_{i1},\ldots,W_{iT})^{\prime}$ is a $T\times d_{\gamma}$ matrix of additional controls associated with common slopes $\gamma_{*}\in\mathbb{R}^{{d_{\gamma}}}$. The common parameters $\gamma_{*}$ can often be identified without knowledge of $G_{*}$ by pooling cross-sectional information.
To see how $\gamma_{*}$ can be identified, rewrite (11) as a pooled-effects model:
$Y_{i}=X_{i}\beta_{*}+W_{i}\gamma_{*}+u_{i},$
where $\beta_{*}:=\mathbb{E}_{G_{*}}[\beta_{i}]$ and $u_{i}=X_{i}(\beta_{i}-\beta_{*})+P_{i}e_{i}$ denotes the composite error. The parameter $\gamma_{*}$ is identified as the familiar regression coefficient
$\gamma_{*}=\mathbb{E}[\tilde{W}_{i}^{\prime}W_{i}]^{-1}\mathbb{E}[\tilde{W}_{i}^{\prime}Y_{i}],$
provided that $\mathbb{E}[\tilde{W}_{i}^{\prime}\tilde{W}_{i}]>0$, where $\tilde{W}_{i}:=W_{i}-X_{i}\mathbb{E}[X_{i}^{\prime}X_{i}]^{-1}\mathbb{E}[X_{i}^{\prime}W_{i}]$. Once $\gamma_{*}$ is identified, the model can be analyzed using the residuals $Y_{i}-W_{i}\gamma_{*}$. Theorem 1 serves as a general tool for analyzing identification in HC models. Before applying Theorem 1 to the HIVDX model, we first consider a HIVD model without the covariate $X_{2i}$. We verify Assumption 2 by demonstrating how to choose a pair $({\boldsymbol{X}},\boldsymbol{M})$.
3.1 Identification of the HIVD Model
We consider two variants of the HIVD model: one with AR(1) errors and the other with ARMA(1,1) errors.
3.1.1 AR(1) Errors
Consider the following AR(1) panel model without covariates:
$Y_{it}$ $=a_{i}+u_{it},$ (12)
$u_{it}$ $=\rho_{i}u_{it-1}+\sigma_{i}e_{it},\quad t=1,\ldots,T,$
where the initial condition $u_{i0}$ is assumed to be drawn from the stationary distribution $\mathcal{N}(0,\sigma_{i}^{2}/(1-\rho_{i}^{2}))$. The parameters are collected in $\theta_{i}=(a_{i},\sigma_{i}^{2},\rho_{i})\overset{\mathrm{i.i.d.}}{\sim}G_{*}$.
Recall the form of the covariance matrix $\Sigma_{i}$ given in (3). This yields the likelihood function evaluated at a generic parameter $\theta=(a,\sigma^{2},\rho)$:
$\ell(Y_{i}\mid\theta)=\frac{(1-\rho^{2})^{1/2}}{(2\pi\sigma^{2})^{T/2}}\exp\left(-\frac{(1-\rho^{2})(Y_{i1}-a)^{2}}{2\sigma^{2}}-\sum_{t=2}^{T}\frac{(Y_{it}-\rho Y_{it-1}-(1-\rho)a)^{2}}{2\sigma^{2}}\right).$ (13)
We establish the identification of $G_{*}$ in the following proposition.
Proposition 1
Assume that $T\geq 3$. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
The restriction on $\mathcal{K}_{(\sigma^{2},\rho)}$ implies that $\rho_{i}$ is bounded away from $\{-1,1\}$ and that $\sigma_{i}^{2}$ is bounded away from 0 and infinity. The condition $T\geq 3$ is not only sufficient but also necessary for the identification of the three-dimensional parameter $\theta_{i}$. The variance parameters are identified from the covariance structure after purging the fixed effect $a_{i}$ via first-differencing $Y_{i}$. Finally, the compact support for $\delta_{i}$ ensures that $Y_{i}$ has non-degenerate and bounded covariance matrices across all $\delta_{i}$; in particular, this condition rules out unit root processes.
We conclude this subsection by illustrating the choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in this example. For simplicity, consider $T=3$. Let $\boldsymbol{X}$} be the column vector of ones whose rank is $1$. Set $\boldsymbol{M}$ equal to the first-difference matrix:
$\boldsymbol{M}=\begin{pmatrix}-1&0\\ 1&-1\\ 0&1\end{pmatrix}$.
This choice of $\boldsymbol{M}$ yields
$\mathcal{V}:=\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}=\frac{\sigma^{2}}{1+\rho}\begin{pmatrix}2&-1+\rho\\ -1+\rho&2\end{pmatrix}$.
Since $\rho=1+2\mathcal{V}_{21}/\mathcal{V}_{11}$ and $\sigma^{2}=\mathcal{V}_{11}+\mathcal{V}_{21}$, there exists a one-to-one relationship between $\mathcal{V}$ and $\delta$, which ensures that $\delta$ is identified from $\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$.
3.1.2 ARMA(1,1) Errors
Consider the ARMA(1,1) panel model:
$Y_{it}$ $=a_{i}+u_{it},$
$u_{it}$ $=\rho_{i}u_{it-1}+\sigma_{i}(e_{it}+\varphi_{i}e_{it-1}),\quad t=1,\dots,T,$
where $e_{it}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1)$ and the initial condition $u_{i0}$ is drawn from the stationary distribution. The additional random coefficient $\varphi_{i}\in[-1,1]$ captures heterogeneity in the MA component. Under the given assumptions, $u_{i}$ is normally distributed with a conditional covariance matrix $\Sigma_{i}$, whose $(t,s)$ entry is given by
$\Sigma_{i,ts}=\frac{\sigma_{i}^{2}}{1-\rho_{i}^{2}}\times\begin{cases}1+\varphi_{i}^{2}+2\varphi_{i}\rho_{i}&\text{if }t=s,\\ (\rho_{i}+\varphi_{i})(1+\rho_{i}\varphi_{i})&\text{if }|t-s|=1,\\ \rho_{i}^{|t-s|-1}(\rho_{i}+\varphi_{i})(1+\rho_{i}\varphi_{i})&\text{if }|t-s|\geq 2.\end{cases}$
Let $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ denote that the probability that $\rho_{i}+\varphi_{i}=0$ equals zero under $G_{*}$.
Proposition 2
Let $\theta_{i}=(a_{i},\sigma_{i}^{2},\rho_{i},\varphi_{i})$ collect all heterogeneous coefficients with $\theta_{i}\overset{\mathrm{i.i.d.}}{\sim}G_{*}$. Assume that $T\geq 4$ and $G_{*}(\rho_{i}+\varphi_{i}=0)=0$. Then $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ and $\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ is a compact subset of $(0,\infty)\times(-1,1)\times[-1,1]$.
The condition $T\geq 4$ is necessary to observe the process over a horizon sufficient to distinguish the two sources of persistence, namely $\rho_{i}$ and $\varphi_{i}$. The restriction $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ ensures that the two dynamic components do not perfectly offset one another. We provide the specific choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in Appendix B.1.
3.2 Identification of the HIVDX Model
We now return to the HIVDX model given in (1). In this model, the parameters are partitioned as $\beta_{i}=(a_{i},b_{i})$ and $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$, yielding the full parameter vector $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ with dimension ${d_{\theta}}=4$. Define $\Delta x_{t}:=x_{t}-x_{t-1}$ as the first difference operator applied to the sequence $(x_{t})_{t=1}^{T}$. Whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariates. We provide two separate identification results: one for $T=4$ and another for $T\geq 5$, which require distinct assumptions on the covariate $X_{2,it}$.
3.2.1 Case I
Suppose that $X_{2,it}$ satisfies either $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0$ or $\Delta X_{2,i2}\neq\Delta X_{2,i4}$ with positive probability. For example, $X_{2,it}=X_{2,i1}+(t-1)$ is a linear trend with a random initial value. Then, we have $\Delta X_{2,it}\equiv 1$ for all $t$.
Proposition 3
Assume that $T=4$ and either $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0\text{ or }\Delta X_{2,i2}\neq\Delta X_{2,i4}$ holds with positive probability. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
To illustrate, consider the case when $\Delta X_{2,it}\equiv 1$. We need to find an appropriate $({\boldsymbol{X}},\boldsymbol{M})$ Pick $X_{2,1}$ in $\operatorname{supp}(X_{2,i1})$ and define
${\boldsymbol{X}}:=\left(\begin{matrix}1&X_{2,1}\\ 1&X_{2,1}+1\\ 1&X_{2,1}+2\\ 1&X_{2,1}+3\end{matrix}\right),\quad\boldsymbol{M}:=\left(\begin{matrix}1&0\\ -2&1\\ 1&-2\\ 0&1\end{matrix}\right),$
where $\boldsymbol{M}$ represents the second-difference operator in matrix form. It is clear that $\boldsymbol{X}$} has full column rank. Furthermore, because $\boldsymbol{M}$ annihilates both the intercept (column of ones) and $X_{2,it}$, the transformation $\boldsymbol{M}^{\prime}Y_{i}$ eliminates the fixed effects $\beta_{i}$ entirely. Let
$\mathcal{V}:=\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}\equiv\mathbb{V}(\boldsymbol{M}^{\prime}Y_{i}\mid\delta_{i}=\delta,X_{i}={\boldsymbol{X}}),$
where $\boldsymbol{M}^{\prime}Y_{i}\in\mathbb{R}^{2}$ collects the second differences of the errors, $(\Delta^{2}{u}_{i3},\Delta^{2}{u}_{i4})$, conditional on $X_{i}={\boldsymbol{X}}$. We obtain the following equations:
$\mathcal{V}_{11}$ $=\mathbb{V}(\Delta^{2}{u}_{i3}\mid\delta_{i}=\delta)=\mathbb{V}(u_{i3}-2u_{i2}+u_{i1}\mid\delta_{i}=\delta)=\frac{2(3-\rho)}{1+\rho}\sigma^{2},$
$\mathcal{V}_{12}$ $=\operatorname{Cov}(\Delta^{2}{u}_{i3},\Delta^{2}{u}_{i4}\mid\delta_{i}=\delta)=\operatorname{Cov}(u_{i3}-2u_{i2}+u_{i1},u_{i4}-2u_{i3}+u_{i2}\mid\delta_{i}=\delta)$
$=\frac{(3-\rho)(\rho-1)}{1+\rho}\sigma^{2}.$
Hence, the system can be inverted:
$(\sigma^{2},\rho)=\left(\frac{\mathcal{V}_{11}}{2}\frac{1+\mathcal{V}_{12}/\mathcal{V}_{11}}{1-\mathcal{V}_{12}/\mathcal{V}_{11}},\frac{2\mathcal{V}_{12}}{\mathcal{V}_{11}}+1\right).$
This shows that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one, thereby establishing Assumption 2(ii).
3.2.2 Case II
Suppose now that $\Delta X_{2,it}$ is not identically zero for some $t\in\{2,\ldots,T\}$. For instance, consider the case where $X_{2,it}$ follows a one-time universal level shift: $X_{2,it}=\mathbf{1}\{t\geq 3\}$, such as a nationwide policy change. Appendix B.1 shows why this pattern fails to satisfy Proposition 3. However, the model is identified, provided $T\geq 5$, as we now show.
Proposition 4
Assume that $T\geq 5$ and that $\Delta X_{2,it}$ is not identically zero for some $t\in\{2,\ldots,T\}$. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
To illustrate, suppose that $X_{2,it}=\mathbf{1}\{t\geq 3\}$. Then, the sequence of first differences is
$(\Delta X_{2,i2},\Delta X_{2,i3},\Delta X_{2,i4},\Delta X_{2,i5})=(0,1,0,0).$
Then, $\boldsymbol{M}$ can now be constructed as
$\boldsymbol{M}=\begin{pmatrix}-1&0&0\\ 1&0&0\\ 0&-1&0\\ 0&1&-1\\ 0&0&1\end{pmatrix}$,
which yields $\boldsymbol{M}^{\prime}Y_{i}=(\Delta u_{i2},\Delta u_{i4},\Delta u_{i5})^{\prime}$. Since the $(1,1)$ and $(2,3)$ elements of the covariance matrix $\mathcal{V}=\mathbb{V}(\boldsymbol{M}^{\prime}Y_{i}\mid\delta_{i}=\delta)$ are given by
$\mathcal{V}_{11}$ $=\mathbb{V}(\Delta{u}_{i2}\mid\delta_{i}=\delta)=\frac{2\sigma^{2}}{1+\rho},$
$\mathcal{V}_{23}$ $=\operatorname{Cov}(\Delta u_{i4},\Delta u_{i5}\mid\delta_{i}=\delta)=\frac{(\rho-1)\sigma^{2}}{1+\rho},$
we can identify $\rho$ via the ratio $2\mathcal{V}_{23}/\mathcal{V}_{11}+1$, and subsequently recover $\sigma^{2}$. This establishes that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one, as required by Assumption 2(ii). By Proposition 4, this result extends to the more general specification
$Y_{it}=a_{i}+b_{i}\,\mathbf{1}\{t\geq s\}+u_{it},\quad t=1,\ldots,5,$
for any known switch point $s\geq 2$, provided the data cover both pre- and post-treatment periods.
As two examples illustrate, incorporating covariates necessitates a significantly more involved analysis. This is because the transformation matrix $\boldsymbol{M}$ forces reliance on differenced data, which often obscures the specific autocovariance structure required to identify $\delta_{i}$. It is worth emphasizing that while the sufficient conditions in Proposition 4 are simple to verify, the underlying proof is combinatorially complex. Establishing this general result requires an exhaustive analysis of all distinct patterns of $\Delta X_{2,it}$ to verify the one-to-one mapping between $\delta$ and $\mathcal{V}$ in every possible scenario.
4 Consistency of the NPMLE
This section establishes that $G_{*}$ can be consistently estimated by the NPMLE under Assumption 2 and the additional regularity assumptions stated below.
Assumption 3.
  1. (i)
    $\mathbb{E}[\|Y_{i}\|^{2}]<\infty$ and $\mathbb{E}[\|X_{i}\|^{2}]<\infty$.
  2. (ii)
    $\mathbb{P}(\operatorname{rank}(X_{i})={d_{\beta}})=1$.
Assumption 3(i) imposes standard moment conditions. Assumption 3(ii) requires full column rank of each $X_{i}$. This ensures that each individual likelihood distinguishes between different values of $\beta_{i}$, so that the individual MLEs are well-defined.
Our proof of the consistency of the NPMLE builds upon Kiefer and Wolfowitz (1956), who consider the case where ${d_{\beta}}=1$. Since we are interested in cases where ${d_{\beta}}>1$, we verify a multivariate version of Assumptions 1–5 in Kiefer and Wolfowitz (1956). Distinct from the issue of identification, the multivariate setting presents a topological complexity not found in the scalar case: probability mass can diverge along multiple directions. A primary technical challenge therefore lies in finding a metric that renders the space $\mathcal{G}$ totally bounded (thereby ensuring the compactness of its completion) while enabling the comparison of distributions with such diverging mass.
To this end, the vague topology is the natural choice. We adopt the canonical metrization of this topology, defined by
$d(G_{0},G_{1})=\sum_{r=1}^{\infty}\frac{1}{2^{r}}\left|\int_{\Theta}h_{r}(\theta)dG_{0}(\theta)-\int_{\Theta}h_{r}(\theta)dG_{1}(\theta)\right|,$
where $\{h_{r}\}_{r=1}^{\infty}$ is a dense sequence in the unit ball of $C_{c}(\Theta)$ with respect to the supremum norm $\|h\|_{\infty}=\sup_{\theta\in\Theta}|h(\theta)|$. [^6] Here, $C_{c}(\Theta)$ denotes the space of continuous real-valued functions on $\Theta$ with compact support. Since the test functions $h_{r}$ vanish at infinity, this metric treats any mass that drifts to infinity as if it were collapsed to a single “point at infinity,” irrespective of the direction of divergence. This choice of metric is compatible with the model because the likelihood cannot distinguish between different parameter values at infinity (i.e., the likelihood vanishes as parameters diverge). In effect, the metric compactifies the domain by collapsing all unbounded directions into a single point, mirroring the likelihood’s asymptotic behavior. Finally, we note that convergence in the vague topology coincides with standard weak convergence whenever the limiting distribution is itself a probability measure.
The next theorem presents the almost sure consistency of the NPMLE with respect to the weak metric $d$.
Theorem 2
Let Assumptions 1, 2, and 3 hold. Then, $d(\hat{G},G_{*})\overset{\operatorname{a.s.}}{\longrightarrow}0$ as $N\to\infty$.
Alternative metrics are discussed in the NPMLE literature, though they serve theoretical roles distinct from our metric $d$. The first is the Hellinger distance between the marginal densities, $\frac{1}{2}\int(\sqrt{f_{\hat{G}}(y)}-\sqrt{f_{G_{*}}(y)})^{2}dy$. While Hellinger distance is frequently considered in mixture models (Jiang and Zhang, 2009; Soloff, Guntuboyina, and Sen, 2025), it measures proximity in the data space rather than the parameter space. In our setting, the convergence of the fitted marginal density $f_{\hat{G}}$ to the truth $f_{G_{*}}$ is of secondary importance. Another alternative is the Wasserstein distance (e.g., $W_{2}$ as in Soloff, Guntuboyina, and Sen, 2025). However, $W_{2}$ requires the convergence of second moments and consequently becomes unbounded if probability mass escapes to infinity. This renders it unsuitable for our initial goal of establishing consistency through compactification. By contrast, our metric $d$ allows us to successfully apply the Kiefer-Wolfowitz machinery. It strikes a necessary balance: it induces a topology sufficiently weak to ensure compactness (accommodating mass at infinity), yet strong enough to imply the consistency of downstream EB estimation.
The consistency of the NPMLE extends to the constrained NPMLE, defined as
$\tilde{G}\in\operatornamewithlimits{arg\hskip 1.00006ptmax}_{G\in\tilde{\mathcal{G}}}F_{N}(G),$
where $\tilde{\mathcal{G}}\subseteq\mathcal{G}$ is a subset that contains $G_{*}$. For instance, Gu and Koenker (2017) employ a profile likelihood approach to estimate a dynamic panel model similar to (12) under the restriction $\rho_{i}=\rho_{*}$. In our framework, imposing homogeneity ($\rho_{i}\equiv\rho$) corresponds to optimizing over a restricted subspace $\tilde{\mathcal{G}}$.
The estimator $\hat{G}$ allows for the recovery of various features of the data-generating process that can be expressed as continuous functionals. In particular, consistent estimators of prior and posterior moments can be obtained from the NPMLE under suitable regularity conditions. For example, the covariance matrix $\mathbb{V}_{G_{*}}(\theta_{i})$ of the true prior can be estimated by $\mathbb{V}_{\hat{G}}(\theta_{i})$, provided that $\|\theta\|^{2}$ is uniformly integrable with respect to the sequence of estimated priors. See the discussion following Assumption 4 for sufficient conditions ensuring such uniform integrability.
5 Regret Consistency and $\ell_{p}$ Convergence
Let $(\tau_{i})_{i=1}^{N}=(\tau(\theta_{i}))_{i=1}^{N}$ denote the unit-level latent parameters of interest, each defined as a function of $\theta_{i}$. [^7] We focus on the case of a scalar parameter here. However, the discussion and results in this section readily extend to the vector-valued case. The estimation problem consists of constructing a decision rule $(Y_{i},X_{i})_{i=1}^{N}\mapsto(\hat{\tau}_{i})_{i=1}^{N}$. Under the compound decision framework, the econometrician seeks to minimize the compound squared-error loss, defined as
$L(\hat{\tau},\tau):=\frac{1}{N}\sum_{i=1}^{N}(\hat{\tau}_{i}-\tau_{i})^{2}.$
The compound risk $R(\hat{\tau},\tau):=\mathbb{E}_{G_{*}}[L(\hat{\tau},\tau)]$ is obtained by taking the expectation of the compound loss jointly over $(Y_{i},X_{i})_{i=1}^{N}$ and $(\theta_{i})_{i=1}^{N}$, accounting for the dependence of the estimator $(\hat{\tau}_{i})_{i=1}^{N}$ on the data. It is well-known that the squared-error risk is minimized by the oracle posterior mean:
${\tau}^{*}_{i}:=\mathbb{E}_{G_{*}}[\tau_{i}\mid(Y_{i},X_{i})_{i=1}^{N}]=\mathbb{E}_{G_{*}}[\tau_{i}\mid Y_{i},X_{i}],\quad i=1,\ldots,N,$ (14)
where the second equality follows from the independence across units. In contrast to the separate decision rule $\hat{\tau}^{\mathrm{MLE}}_{i}=\tau(\hat{\theta}^{\operatorname{MLE}}_{i})$, efficiency gains are achieved by pooling information across units through $G_{*}$. G-modeling approximates the oracle (14) by substituting $\hat{G}$ for $G_{*}$, yielding the estimator
$\hat{\tau}^{\mathrm{EB}}_{i}:=\mathbb{E}_{\hat{G}}[\tau_{i}\mid Y_{i},X_{i}]\equiv\frac{\int_{\Theta}\tau(\theta)\ell(Y_{i}\mid X_{i},\theta)d\hat{G}(\theta)}{f_{\hat{G}}(Y_{i},X_{i})}.$
This section establishes asymptotic guarantees for the EB method in terms of regret, defined as:
$R(\hat{\tau}^{\mathrm{EB}},\tau)-R({\tau}^{*},\tau)=\mathbb{E}_{G_{*}}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{\tau}^{\mathrm{EB}}_{i}-\tau_{i})^{2}-({\tau}^{*}_{i}-\tau_{i})^{2}\right]=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i})^{2}\right]\geq 0,$
since $\mathbb{E}_{G_{*}}[(\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i})({\tau}^{*}_{i}-\tau_{i})]=0$ by the properties of conditional expectation. The condition $\lim_{N\to\infty}[R(\hat{\tau}^{\mathrm{EB}},\tau)-R({\tau}^{*},\tau)]=0$ is referred to as regret consistency (see, e.g., Adusumilli, Gu, and Tao (2025)) or asymptotic optimality (see, e.g., Zhang (1997)). We mainly refer to this condition as regret consistency.
Consider the case when $\tau_{i}$ is the prediction of $Y_{iT+1}$. The oracle predictor ${\tau}^{*}_{i}$ in the HIVDX model is given by
${\tau}^{*}_{i}$ $=\mathbb{E}_{G_{*}}[a_{i}+b_{i}X_{2,iT+1}+\rho_{i}Y_{iT}\mid Y_{i},X_{i},X_{2,iT+1}]-\mathbb{E}_{G_{*}}[\rho_{i}a_{i}+\rho_{i}b_{i}X_{2,iT}\mid Y_{i},X_{i}].$ (15)
The EB one-period-ahead prediction $\hat{\tau}^{\mathrm{EB}}_{i}$ is then obtained by substituting the NPMLE for $G_{*}$ in (15). Evaluating the EB method in this setting requires controlling the compound error arising from the interaction between the estimation discrepancy $\hat{\theta}^{\mathrm{EB}}_{i}-{\theta}^{*}_{i}$ and these observed variables. Consequently, the analysis necessitates controlling the $\ell_{p}$ distance (where $p\geq 2$, with $p=4$ being a particularly important case) between the EB and oracle estimates. We now establish regret consistency under the high-level assumptions stated below.
Assumption 4.
Assume that $\mathbb{E}_{G_{*}}[|\tau_{i}|^{p}]<\infty$ for some $p\in[2,\infty)$. Moreover, the following conditions hold.
  1. (i)
    The function $\tau(\theta)$ is continuous in $\theta\in\Theta$ and, for all $c>0$, satisfies
    $\lim_{\|\beta\|\to\infty}\exp(-c\|\beta\|^{2})|\tau(\theta)|=0.$
  2. (ii)
    For all $\epsilon>0$, there exists $M<\infty$ such that
    $\limsup_{N\to\infty}\mathbb{E}\left[\int_{\Theta}|\tau(\theta)|^{p}\mathbf{1}\{|\tau(\theta)|\geq M\}d\hat{G}_{N}(\theta)\right]$ $\leq\epsilon,$
    where $\hat{G}_{N}$ is the NPMLE defined in (7), with NN denoting the sample size.
Assumption 4(i) requires that $\tau(\theta)$ depends continuously on $\theta$ and grows slower than any quadratic exponential function of $\|\beta\|$. This ensures that $|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|\to 0$ for each unit $i$ as $d(\hat{G}_{N},G_{*})\to 0$. The continuity requirement can be relaxed to allow $\tau(\theta)$ to be discontinuous on a set where $G_{*}$ places zero mass.
Assumption 4(ii) imposes a high-level condition on the NPMLE sequence, requiring $|\tau_{i}|^{p}$ to be uniformly integrable with respect to the sequence of estimated priors. This condition is automatically satisfied when $\tau_{i}$ has bounded support (e.g., $\tau_{i}=\delta_{i}$) but can be nontrivial for unbounded $\tau_{i}$ such as slope parameters. For such cases, Assumption 4(ii) requires specific moment conditions on $(Y_{i},X_{i})$; for example, the $p$-th moment of $\beta_{i}$ in the HIVDX model is bounded by the $p$-th moment of $Y_{i}$, provided $X_{i}$ satisfies certain regularity conditions.
Theorem 3
Let Assumptions 1, 2, 3, and 4 hold for some $p\in[2,\infty)$. Then $(\hat{\tau}^{\mathrm{EB}}_{i})_{i=1}^{N}$ is regret-consistent, i.e.,
[…]
Moreover, the EB estimators converge in $\ell_{p}$-norm:
$\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{p}\right]\to 0\quad\text{as }N\to\infty.$
Theorem 3 provides the theoretical justification for using the NPMLE-based EB rule in large samples. The primary challenge in establishing this result is that the weak convergence of the priors, $d(\hat{G},G_{*})\overset{\operatorname{a.s.}}{\longrightarrow}0$, implies the convergence of integrals only for bounded continuous functions. However, $\tau(\theta)$ and the squared-error loss are generally unbounded functions of the parameters. Assumption 4(ii) is critical because uniform integrability of the estimated prior moments ensures uniform integrability of the posterior moments across units. This allows us to upgrade the weak convergence of $\hat{G}$ to the convergence of the posterior expectations (the EB estimators) in $\ell_{p}$-norm. Notably, the second result of the theorem, that is, convergence in $\ell_{p}$-norm, is stronger than simple regret consistency ($p=2$) and is particularly useful when the latent parameters enter the decision problem nonlinearly or interact with unbounded covariates.
5.1 Regret Consistency in the HIVDX Model
This subsection applies Theorem 3 to the HIVDX model, focusing on several quantities of interest: the individual-level parameters $\theta_{i}$ and the one-period-ahead prediction. These examples provide the theoretical underpinnings for the empirical analysis presented in Section 7.
When $\tau_{i}=\theta_{i}$, the EB estimator of $\theta_{i}$ is given by
$\hat{\theta}^{\mathrm{EB}}_{i}:=\frac{\int_{\Theta}\theta\ell(Y_{i}\mid X_{i},\theta)d\hat{G}(\theta)}{f_{\hat{G}}(Y_{i},X_{i})},$
which can be partitioned as $\hat{\theta}^{\mathrm{EB}}_{i}=(\hat{\beta}^{\mathrm{EB}}_{i},\hat{\delta}^{\mathrm{EB}}_{i})$. The regret consistency of $\hat{\delta}^{\mathrm{EB}}_{i}$ follows immediately from Theorem 3 and Assumption 4, as $\delta_{i}$ lies in a compact set. For $\hat{\beta}^{\mathrm{EB}}_{i}$, however, Assumption 4(ii) requires $\|\beta_{i}\|^{p}$ to be uniformly integrable with respect to the sequence of estimated priors. This condition is guaranteed by a moment condition on $(Y_{i},X_{i})$.
Once $(\theta_{i})_{i=1}^{N}$ have been estimated, we turn to predicting future outcomes. We focus on the one-period-ahead EB prediction $\hat{Y}^{\mathrm{EB}}_{iT+1}$ of $Y_{iT+1}$ (recall the oracle estimator given in (15)). Let $\bar{X}_{2,i}=T^{-1}\sum_{t=1}^{T}X_{2,it}$ denote the time average of $X_{2,it}$ for unit $i$.
Proposition 5
Let Assumption 1 hold. Assume that $|\bar{X}_{2,i}|\leq M$ and $\sum_{t=1}^{T}(X_{2,it}-\bar{X}_{2,i})^{2}\geq c$ for some constants $M,c>0$ and for all $i\geq 1$. Then, we have the following regret consistency results:
  1. (i)
    $(\hat{\sigma}^{2,\mathrm{EB}}_{i})_{i=1}^{N}$ and $(\hat{\rho}^{\mathrm{EB}}_{i})_{i=1}^{N}$ are regret-consistent.
  2. (ii)
    In addition, assume that $\mathbb{E}[\|Y_{i}\|^{2+\varepsilon}]<\infty$ for some $\varepsilon>0$. Then $(\hat{a}^{{\mathrm{EB}}}_{i})_{i=1}^{N}$ and $(\hat{b}^{{\mathrm{EB}}}_{i})_{i=1}^{N}$ are regret-consistent.
  3. (iii)
    Assume further that $\mathbb{E}[\|Y_{i}\|^{4+\varepsilon}]<\infty$ for some $\varepsilon>0$, $\mathbb{E}[|X_{2,iT}|^{4}]<\infty$, and $\mathbb{E}[|X_{2,iT+1}|^{4}]<\infty$. Then the predictions $(\hat{Y}^{{\mathrm{EB}}}_{iT+1})_{i=1}^{N}$ are regret-consistent.
Proposition 5(iii) requires fourth (and slightly higher) moments of $(X_{2,iT},X_{2,iT+1})$ and $Y_{i}$ to control the components of the prediction risk $\mathbb{E}[N^{-1}\sum_{i=1}^{N}(\hat{Y}^{{\mathrm{EB}}}_{iT+1}-\hat{Y}^{*}_{iT+1})^{2}]$. For instance, applying the Cauchy-Schwarz inequality to the component involving the slope coefficient yields
$\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{b}^{{\mathrm{EB}}}_{i}-\hat{b}^{*}_{i})^{2}X_{2,iT+1}^{2}\right]\leq\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{b}^{{\mathrm{EB}}}_{i}-\hat{b}^{*}_{i})^{4}\right]}\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}X_{2,iT+1}^{4}\right]}.$
The moment assumption on $Y_{i}$ ensures that the first term on the right-hand side vanishes asymptotically (by Theorem 3 with $p=4$), while the fourth moment of $X_{2,iT+1}$ ensures that the second term remains bounded.
6 Computational Algorithms
A variety of computational methods have been proposed for implementing the NPMLE. Although the optimization problem is convex, computation is challenging because the optimization variable $G$ is an infinite-dimensional object, rendering direct implementation intractable. A common remedy is to approximate $G$ by a discrete distribution supported on $m$ atoms (with $m\in\mathbb{N}$), thereby reducing the problem to fitting an $m$-component mixture model. [^8] This approach is justified by the well-known result that a discrete NPMLE with at most $N_{0}$ support points exists (possibly among multiple maximizers), where $N_{0}$ is the number of distinct data points (Lindsay, 1983). Furthermore, Shen and Wu (2025) show that the required number of components $m$ grows at the rate $O(\log N)$ in the one-dimensional case under subgaussian priors, and conjecture that in higher dimensions $m$ increases on the order of $O((\log N)^{C})$, where $C$ depends on the dimension of $\theta_{i}$. For the specific case of one-dimensional Gaussian location mixtures, Polyanskiy and Sellke (2025) establish computational guarantees, including an algorithm that computes the exact support size.
For example, Jiang and Zhang (2009) introduce a fixed-point EM algorithm that mirrors the standard EM routine and establish that it attains a global optimum up to an explicit error bound. In practice, however, EM with a large fixed grid often converges prohibitively slowly, making it unsuitable for our HC model. Koenker and Mizera (2014) develop a convex-optimization method based on an interior-point algorithm to fit weights on a fixed grid, which Gu and Koenker (2017) apply to a dynamic panel model with two-dimensional heterogeneity. While generally reliable, this approach requires a prespecified grid whose size grows exponentially with the dimension of the support. More recently, Kim, Carbonetto, Stephens, and Anitescu (2020), Zhang, Cui, Sen, and Toh (2024), and Wang, Ibrahim, and Mazumder (2025) propose alternative methods based on general-purpose convex programming to improve scalability. However, high-dimensional computational challenges persist. For instance, Zhang, Cui, Sen, and Toh (2024) suggest using observed data points as the grid when the dimension of multivariate mixture models is three or higher.
A feature shared by all the methods above is that the location of the support points is fixed; optimization is performed solely with respect to the weights. In our HIVDX model ($d_{\theta}=4$), this imposes a severe trade-off: a grid dense enough to minimize approximation error becomes computationally intractable due to the curse of dimensionality, while a sparse grid risks significant discretization bias. Moreover, unlike in simple mixture models, there are no natural candidates (such as observed data points) to serve as a baseline grid in our setting. Consequently, these recent proposals are not well-suited for our purposes. Motivated by these challenges for multiple sources of heterogeneity, we consider a discrete-time Wasserstein-Fisher-Rao (WFR) gradient flow algorithm on the space of probability measures over $\Theta$. Crucially, the WFR algorithm adaptively relocates support points; as a result, the number of support atoms $m$ remains moderate and need not exceed the sample size $N$. We now explain our procedure in detail.
Let $\boldsymbol{\theta}=(\theta^{j})_{j=1}^{m}$ denote a grid of support points for $G$, and let $\boldsymbol{\omega}=(\omega^{j})_{j=1}^{m}$ be the associated weights satisfying $\omega^{j}\geq 0$ for all $j$ and $\sum_{j=1}^{m}\omega^{j}=1$. We use the superscript $j$ to distinguish the grid $(\theta^{j})_{j=1}^{m}\in\Theta^{m}$ from the individual-specific parameters $(\theta_{i})_{i=1}^{N}\in\Theta^{N}$. Let $G^{\boldsymbol{\theta},\boldsymbol{\omega}}$ denote the discrete distribution induced by $(\boldsymbol{\theta},\boldsymbol{\omega})$, i.e.,
$G^{\boldsymbol{\theta},\boldsymbol{\omega}}(\Theta^{\prime})=\sum_{j=1}^{m}\omega^{j}\,{\mathbf{1}\large[{\theta^{j}\in\Theta^{\prime}}\large]}\qquad\text{for all subsets }\Theta^{\prime}\subseteq\Theta.$
The NPMLE of the HC model can then be expressed as the solution to
$\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m},\,\boldsymbol{\omega}\geq 0}F_{N}(G^{\boldsymbol{\theta},\boldsymbol{\omega}})=\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m},\,\boldsymbol{\omega}\geq 0}\frac{1}{N}\sum_{i=1}^{N}\log f_{G^{\boldsymbol{\theta},\boldsymbol{\omega}}}(Y_{i},X_{i})\quad\text{s.t. }\sum_{j=1}^{m}\omega^{j}=1,$
where
$f_{G^{\boldsymbol{\theta},\boldsymbol{\omega}}}(Y_{i},X_{i})=\int_{\Theta}\ell(Y_{i}\,|\,X_{i},\theta)\,dG^{\boldsymbol{\theta},\boldsymbol{\omega}}(\theta)=\sum_{j=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{j})\,\omega^{j}.$
By the Karush-Kuhn-Tucker theorem, for some $\lambda\geq 0$, this problem is equivalent to
$\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m},\,\boldsymbol{\omega}\geq 0}\frac{1}{N}\sum_{i=1}^{N}\log\left(\sum_{j=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{j})\omega^{j}\right)-\lambda\sum_{j=1}^{m}\omega^{j},$ (16)
where the value of $\lambda=1$ can be determined by investigating the first-order conditions for each weight $\omega^{j}$.
Even with discretization, finding a path to the global maximizer remains challenging because the space of probability measures denoted by $\mathcal{P}(\Theta)$ has a complex geometry, and different metrics can generate markedly different gradient-flow trajectories. [^9] Gradient flow generalizes gradient descent to metric spaces beyond the standard Euclidean setting. In our case, the objective is the likelihood function, and the decision variable $G$ lies in the space of probability measures. For example, the commonly used Fisher-Rao (FR) metric and the quadratic Wasserstein (W) distance induce distinct notions of gradient flow on $\mathcal{P}(\Theta)$. The FR gradient flow updates $G_{t}^{\mathrm{FR}}$ according to
$G_{t+\Delta t}^{\mathrm{FR}}-G_{t}^{\mathrm{FR}}\approx G_{t}^{\mathrm{FR}}\big(\alpha_{t}-\mathbb{E}_{G_{t}^{\mathrm{FR}}}[\alpha_{t}]\big),$
where the reweighting strategy $\alpha_{t}:\Theta\to\mathbb{R}$ is chosen to maximize $F_{N}(G_{t+\Delta t}^{\mathrm{FR}})-F_{N}(G_{t}^{\mathrm{FR}})$ subject to the normalization $\mathbb{E}_{G_{t}^{\mathrm{FR}}}[(\alpha_{t}-\mathbb{E}_{G_{t}^{\mathrm{FR}}}[\alpha_{t}])^{2}]=1$. The limiting path $(G_{t}^{\mathrm{FR}})_{t\geq 0}$ as $\Delta t\to 0$ defines the FR gradient flow. In contrast, the W gradient flow evolves according to the continuity (transport) equation
$G_{t+\Delta t}^{\mathrm{W}}-G_{t}^{\mathrm{W}}\approx-\mathrm{div}(G_{t}^{\mathrm{W}}V_{t}),$
where the velocity field $V_{t}:\Theta\to\mathbb{R}^{{d_{\theta}}}$ describes the movement of mass and is chosen to maximize the gain in $F_{N}$ under the normalization $\mathbb{E}_{G_{t}^{\mathrm{W}}}[\|V_{t}\|^{2}]=1$. The resulting trajectory $(G_{t}^{\mathrm{W}})_{t\geq 0}$ is referred to as the W gradient flow. Because they are derived from different metrics, the FR and W flows yield different steady-state conditions. Notably, the first-order condition of the FR flow implies that of the W flow. The FR flow is guaranteed to reach a global maximum when initialized with full support, as established by Yan, Wang, and Rigollet (2024) for the Gaussian location mixture case, whereas the W flow may stall at a local maximum. Intuitively, the FR flow emphasizes reweighting existing support points, while the W flow emphasizes transporting mass across the space, leading to different convergence behavior.
The Wasserstein-Fisher-Rao (WFR) gradient flow combines both reweighting and transport dynamics when updating $G_{t}^{\mathrm{WFR}}$:
$G_{t+\Delta t}^{\mathrm{WFR}}-G_{t}^{\mathrm{WFR}}\approx-\mathrm{div}(G_{t}^{\mathrm{WFR}}V_{t})+G_{t}^{\mathrm{WFR}}\big(\alpha_{t}-\mathbb{E}_{G_{t}^{\mathrm{WFR}}}[\alpha_{t}]\big).$
As a result, the WFR gradient flow inherits the global convergence guarantees of the FR flow, while the W component allows it to adapt the support through mass transport. In the context of Gaussian location mixtures, Yan, Wang, and Rigollet (2024) provide theoretical guarantees for a discrete-time WFR gradient-descent algorithm and empirically demonstrates its superiority over both the EM algorithm and gradient-descent methods based solely on the FR or W flow. We extend their approach to a more general HC model that includes the HIVD and HIVDX models as special cases. Intuitively, WFR balances exploration through transport with refinement through reweighting, which explains its superior performance in both theory and practice. Further details are presented in Section 6.1, with a comparison to the EM algorithm in Section 6.2.
6.1 WFR Algorithm for the HC Model
The WFR gradient-descent algorithm updates $(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n})_{n\geq 0}$ by alternately reweighting and moving the mass along the steepest descent path. Let $\eta\in(0,1]$ denote the step size (learning rate), and define the posterior weights
$\pi_{ij}(\boldsymbol{\theta},\boldsymbol{\omega})=\frac{\ell(Y_{i}\,|\,X_{i},\theta^{j})\,\omega^{j}}{\sum_{k=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{k})\,\omega^{k}},\qquad i=1,\ldots,N,\ j=1,\ldots,m,$
which correspond to the E-step in the EM algorithm. Holding $\boldsymbol{\theta}_{n}$ fixed, the algorithm first updates the weights $\boldsymbol{\omega}_{n}$ via
$\begin{aligned} \omega_{n+1}^{j}&=\omega_{n}^{j}\left(1+\eta\left[\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta_{n}^{j})}{\sum_{k=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta_{n}^{k})\,\omega_{n}^{k}}-1\right]\right)\\ &=(1-\eta)\omega_{n}^{j}+\frac{\eta}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n}),\qquad j=1,\ldots,m,\end{aligned}$ (FR-step)
which we refer to as the FR-step. The reweighting function is chosen as the gradient of the objective function in (16) with respect to $\omega^{j}$, thereby ensuring descent. Next, with $\boldsymbol{\omega}_{n+1}$ fixed, the grid $\boldsymbol{\theta}_{n}$ is updated according to
$\begin{aligned} \tilde{\theta}_{n+1}^{j}&=\theta_{n}^{j}+\frac{\eta}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n+1})\frac{\partial}{\partial\theta_{n}^{j}}\log\ell(Y_{i}\,|\,X_{i},\theta_{n}^{j}),\\ \theta_{n+1}^{j}&=P_{\Theta}(\tilde{\theta}_{n+1}^{j}),\qquad j=1,\ldots,m,\end{aligned}$ (W-step)
[…]
$\max_{j}\left\{\frac{1}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n})-1\right\}\leq\mathtt{tol}$
for some tolerance parameter $\mathtt{tol}>0$. [^11] A necessary and sufficient condition for $\hat{G}$ to be NPMLE is $\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\leq 1\text{ for all }\theta\in\Theta.$ Let $\hat{n}$ denote the stopping index (either the first $n$ satisfying the convergence criterion or $\overline{n}$ if it is not met). The resulting estimate is the NPMLE $\hat{G}=G^{\boldsymbol{\theta}_{\hat{n}},\boldsymbol{\omega}_{\hat{n}}}$. Implementation of the HC model is summarized in Algorithm 1.
Input: Dataset $(Y_{i},X_{i})_{i=1}^{N}$; initial grid and weights $(\theta_{0}^{j},\omega_{0}^{j})_{j=1}^{m}$; step size $\eta\in(0,1]$; maximum iterations $\overline{n}$; tolerance parameter $\mathtt{tol}$.
Set $n\leftarrow 0$.
while $n<\overline{n}\;\;\mathrm{and}\;\max_{j}\{\frac{1}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n})-1\}>\mathtt{tol}$ do
 /* FR-step: reweighting */
 for $j=1,\ldots,m$ do
     Compute posterior weights $\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n})=\frac{\ell(Y_{i}\mid X_{i},\theta_{n}^{j})\,\omega_{n}^{j}}{\sum_{k=1}^{m}\ell(Y_{i}\mid X_{i},\theta_{n}^{k})\,\omega_{n}^{k}},\quad i=1,\ldots,N.$ Update $\omega_{n+1}^{j}=(1-\eta)\,\omega_{n}^{j}+\frac{\eta}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n}).$
   end for
 /* W-step: transport */
 for $j=1,\ldots,m$ do
     Compute updated posteriors (with new weights) $\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n+1})=\frac{\ell(Y_{i}\mid X_{i},\theta_{n}^{j})\,\omega_{n+1}^{j}}{\sum_{k=1}^{m}\ell(Y_{i}\mid X_{i},\theta_{n}^{k})\,\omega_{n+1}^{k}},\quad i=1,\ldots,N.$
     Take a gradient step $\tilde{\theta}_{n+1}^{j}=\theta_{n}^{j}+\frac{\eta}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n+1})\,\frac{\partial}{\partial\theta_{n}^{j}}\log\ell(Y_{i}\,|\,X_{i},\theta_{n}^{j}).$
     Project to the feasible set $\theta_{n+1}^{j}\leftarrow P_{\Theta}(\tilde{\theta}_{n+1}^{j}).$
   end for
  Set $n\leftarrow n+1$.
end while
Output: Final grid and weights $(\theta_{\hat{n}}^{j},\omega_{\hat{n}}^{j})_{j=1}^{m}$, where $\hat{n}$ denotes the stopping index (the first $n$ satisfying the convergence criterion, or $\overline{n}$ if it is not met).
Algorithm 1 Wasserstein-Fisher-Rao algorithm for the HC model
Starting values
Choosing an initial grid that is well spread over $\Theta$ is essential for the WFR gradient-descent algorithm to converge to a global optimum (Yan, Wang, and Rigollet, 2024, Theorem 4). If, instead, all atoms are initialized at the same location, i.e., $\theta_{0}^{j}=\tilde{\theta}$ for all $j=1,\ldots,m$, then the subsequent updates satisfy $\theta_{n}^{1}=\cdots=\theta_{n}^{m}$ for every $n\geq 0$. In this case, the distribution $G^{\boldsymbol{\theta}_{n},\boldsymbol{\omega}_{n}}$ degenerates to a single point mass at $\theta_{n}^{j}=\tilde{\theta}$, failing to capture any heterogeneity in the $\theta_{i}$’s and hence unable to recover the NPMLE.
In practice, the following procedure can be used to initialize the support points and weights. For $j=1,\ldots,m$:
  • (i)
    Sample indices $i(1),\ldots,i(B)$ uniformly at random from $\{1,\ldots,n\}$.
  • (ii)
    Compute
    $\theta_{0}^{j}=\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\theta\in\Theta}\sum_{b=1}^{B}\log\ell(Y_{i(b)}\,|\,X_{i(b)},\theta)$
    as the subsample MLE based on $\{(Y_{i(b)},X_{i(b)})\}_{b=1}^{B}$, and assign a uniform weight $\omega_{0}^{j}=1/m$.
The resulting collection $(\theta_{0}^{j},\omega_{0}^{j})_{j=1}^{m}$ serves as the initial grid and weights.
6.2 Comparison with the EM Algorithm
The EM algorithm is often used to find a (local) solution to mixture problems of the form (16). Let $D_{ij}={\mathbf{1}\large[{\theta_{i}=\theta^{j}}\large]}$ denote an unobserved indicator of unit $i$’s membership in component $j$. If the true $D_{ij}$ were known, the “complete-data” maximum likelihood problem would be
$\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m}}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{m}D_{ij}\log\ell(Y_{i}\,|\,X_{i},\theta^{j})=\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m}}\sum_{j=1}^{m}\sum_{i:D_{ij}=1}\log\ell(Y_{i}\,|\,X_{i},\theta^{j}),$
which can be optimized directly with respect to $\boldsymbol{\theta}$. However, this is infeasible because the true $D_{ij}$ are unobserved. The EM algorithm circumvents this difficulty by replacing $D_{ij}$ with its conditional expectation,
$\mathbb{E}_{\boldsymbol{\theta},\boldsymbol{\omega}}[D_{ij}\,|\,Y_{i},X_{i}]=\frac{\ell(Y_{i}\,|\,X_{i},\theta^{j})\,\omega^{j}}{\sum_{k=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{k})\,\omega^{k}}=\pi_{ij}(\boldsymbol{\theta},\boldsymbol{\omega}),$
evaluated at the current iterate $(\boldsymbol{\theta},\boldsymbol{\omega})$. This step is referred to as the E-step. Let $(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}})$ denote the $n$th iterate of the support points and their weights. Given these values, the expected log-likelihood can be written as
$Q_{N}(\boldsymbol{\theta}\,|\,\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}}):=\frac{1}{N}\sum_{j=1}^{m}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}})\log\ell(Y_{i}\,|\,X_{i},\theta^{j}).$
In the subsequent M-step, $(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}})$ are updated by solving
$\theta_{n+1}^{j,\mathrm{EM}}$ $=\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}})\log\ell(Y_{i}\,|\,X_{i},\theta),$
$\omega_{n+1}^{j,\mathrm{EM}}$ $=\frac{1}{N}\sum_{i=1}^{N}\pi_{ij}(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}}),\qquad j=1,\ldots,m.$
The E- and M-steps are iterated until convergence, $(\boldsymbol{\theta}_{n}^{\mathrm{EM}},\boldsymbol{\omega}_{n}^{\mathrm{EM}})\to(\hat{\boldsymbol{\theta}}^{\mathrm{EM}},\hat{\boldsymbol{\omega}}^{\mathrm{EM}})$, yielding the EM estimator $\hat{G}^{\mathrm{EM}}=G^{\hat{\boldsymbol{\theta}}^{\mathrm{EM}},\hat{\boldsymbol{\omega}}^{\mathrm{EM}}}$.
When $\eta$ in our algorithm is set to $1$, the FR-step coincides with the weight update in the EM algorithm. For smaller values of $\eta$, the weights are adjusted more conservatively, leading to more stable updates than in EM. Similarly, the W-step can be viewed as moving $\boldsymbol{\theta}_{n}$ incrementally along the gradient of the $Q_{N}$-function. Because it is formulated as a gradient-descent procedure, the WFR algorithm enjoys stronger convergence guarantees than EM. By contrast, the EM algorithm generally requires more careful initialization to reach the global maximum (Balakrishnan, Wainwright, and Yu, 2017).
7 Empirical Analysis of Income Dynamics
There is a large body of literature seeking to understand heterogeneity in earnings. See Altonji, Hynsjö, and Vidangos (2023) and Browning and Ejrnaes (2013) for recent reviews of this literature. We highlight several closely related methodological contributions. Chamberlain and Hirano (1999) consider a parametric Bayesian approach that allow for heterogeneity in $\sigma_{i}^{2}$. Geweke and Keane (2000) also adopt a Bayesian framework with a finite mixture model for the composite errors $\sigma_{i}e_{it}$, while Hirano (2002) extend this framework using an infinite mixture model. Gu and Koenker (2017) propose an EB approach that was closest in spirit to ours. These four papers did not allow for random $(b_{i},\rho_{i})$. Giacomini, Lee, and Sarpietro (2025) develop an individual-weight shrinkage estimator that optimizes unit-level (rather than aggregate) accuracy by exploiting each individual’s own past history, with feasible weights motivated by minimax regret. Moon, Schorfheide, and Zhang (2023) consider random components in all $(a_{i},b_{i},\rho_{i},\sigma_{i}^{2})$ within a parametric Bayesian framework. Specifically, they adopt a spike-and-slab prior that accommodate either dense or sparse structures but assume independence across random components. In contrast, we adopt a nonparametric EB approach for the prior distribution and allow for dependence among the random components.
Gu and Koenker (2017, GK hereafter) analyze log real earnings data studied in Meghir and Pistaferri (2004). The extract consists of 938 individuals who have continuous earnings records from age 25 onward in the Panel Study of Income Dynamics (PSID) for the period 1968-1993. The panel is unbalanced panels with varying numbers of time periods $T_{i}$. [^12] The only minor difference between the current model and the original HIVDX model in (1) is that $T_{i}$ may vary across $i$. The likelihood function can be easily modified to accommodate unbalanced panels. Their $Y_{it}$ denotes residuals obtained from year-specific regressions of log real earnings on a vector of covariates. Though the HIVDX model motivates GK’s analysis, they impose $\rho_{i}=\rho_{*}$ for some unknown homogeneous parameter $\rho_{*}$ and $b_{i}\equiv 0$ to circumvent computational challenges. Furthermore, their likelihood is conditioned on the initial observation $Y_{i1}$, while implicitly assuming that the random component is independent of $Y_{i1}$. Thus, the marginal likelihood depends on $Y_{i1}$ only through $Y_{i2}-\rho_{*}Y_{i1}$.
Using GK’s PSID sample extract and their definition of $Y_{it}$ (residualized incomes), we estimate the HIVDX model with $\theta_{i}:=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ and $X_{2,it}:=\mathrm{Exp}_{it}/10$, so that $b_{i}$ is scaled up by a factor of $10$, and $X_{2,it}$ is potential experience constructed as $\mathrm{Exp}_{it}\;:=\;\text{age}\;-\;\max\{\text{years of schooling},\,12\}\;-\;6$. [^13] An equivalent form used in the literature is $\mathrm{Exp}_{it}=\min\{\text{age}-\text{years of schooling}-6,\text{age}-18\}$. Furthermore, our analysis differs from GK in three other respects. First, we model $Y_{i1}$ under a stationarity assumption without assuming independence between $Y_{i1}$ and $\theta_{i}$. Second, the likelihood is directly evaluated from the observed data since sufficient statistics are not available in the presence of $X_{2,it}$ and $\rho_{i}$. Third, we fit the model and estimate the prior distribution of $\theta_{i}$ using the WFR algorithm described in Section 6.1. In empirical analyses, we used $\overline{n}=100,000$ as the maximum number of iterations and did not implement the stopping criterion. The initialization algorithm with $B=1$ and $m=500$ and the step size of $\eta=0.0005$ are used.
Table 1: First and Second Moments of the Estimated Prior $\hat{G}$
$\mathbb{V}_{\hat{G}}(\theta_{i})$ $\theta_{i}$
$a_{i}$ $b_{i}$ $\sigma_{i}^{2}$ $\rho_{i}$
$a_{i}$ $0.2052$ 0.1003-0.1003 0.0105-0.0105 0.0039-0.0039
$b_{i}$ 0.1003-0.1003 $0.1007$ 0.0104-0.0104 $0.0074$
$\sigma_{i}^{2}$ 0.0105-0.0105 0.0104-0.0104 $0.0229$ 0.0013-0.0013
$\rho_{i}$ 0.0039-0.0039 $0.0074$ 0.0013-0.0013 $0.1052$
$\mathbb{E}_{\hat{G}}[\theta_{i}]$ $0.0134$ 0.0166-0.0166 $0.0787$ $0.4219$
Notes: The first four rows form a $4\times 4$ covariance matrix. The last row presents means.
Table 1 shows the first and second moments from the estimated prior $\hat{G}$. The marginal variances suggest that the coefficients are heterogeneous in all four dimensions. The variance of $a_{i}$ is substantial, especially relative to the variance of $Y_{it}$ across all units and time ($0.22$). Regarding $b_{i}$, the variance of $b_{i}X_{2,it}$ when $X_{2,it}=x$ is approximately $0.1x^{2}$. The value of $x$ ranges over $[0.2,3.2]$ within the sample, indicating notable heterogeneity in experience profiles. The variance of $\sigma_{i}^{2}$ is 0.0229 and is non-negligible. Finally, the distribution of $\rho_{i}$ is also spread out, with approximately $73\%$ of probability mass lying in $[0.2,0.8]$.
Turning to the covariances of the random coefficients, the covariance between individual intercepts $(a_{i})$ and variances $(\sigma_{i}^{2})$ is negative, as in GK. However, allowing for additional sources of heterogeneity leads to some new findings. We find a negative correlation between $\sigma_{i}^{2}$ and $b_{i}$, and weak associations between $\rho_{i}$ with the other parameters ($a_{i},b_{i},\sigma_{i}^{2}$). The most interesting result is the strong negative correlation between the individual intercepts ($a_{i}$) and slopes ($b_{i}$).
Figure 1: Shrinkage Effects: Comparison of $(\hat{\theta}^{\mathrm{MLE}}_{i})_{i=1}^{N}$ and $(\hat{\theta}^{\mathrm{EB}}_{i})_{i=1}^{N}$
(a) MLEs of $a_{i}$.
(b) EB estimates of $a_{i}$.
(c) MLEs of $b_{i}$.
(d) EB estimates of $b_{i}$.
(e) MLEs of $\rho_{i}$.
(f) EB estimates of $\rho_{i}$.
(g) MLEs of $\sigma_{i}^{2}$.
(h) EB estimates of $\sigma_{i}^{2}$.
Note: The left and right panels display the histograms of the individual MLE and EB estimates for each component of $(\theta_{i})_{i=1}^{N}$.
A side-by-side comparison between the individual MLEs and EB estimates of $\theta_{i}$ is presented in Figure 1. As we work with residual earnings, it is not surprising that the means of the individual MLEs and EB estimates of $a_{i}$ and $b_{i}$ are close to zero. A notable feature is that the individual MLEs of $\rho_{i}$ are highly dispersed, with a substantial fraction taking negative values, whereas the EB estimates are mostly positive and considerably more concentrated around their mean value of 0.422. The cross-sectional variances of $\hat{a}_{i}^{\mathrm{EB}}$, $\hat{b}_{i}^{\mathrm{EB}}$, and $\hat{\rho}_{i}^{\mathrm{EB}}$ are reduced by approximately 63%, 72%, and 48%, respectively, relative to the corresponding individual MLEs. While no conspicuous shrinkage is observed for $\hat{\sigma}_{i}^{2,\mathrm{EB}}$, these reductions suggest substantial compound mean squared error gains from the EB approach.
Lastly, we consider one-period-ahead predictions of $(Y_{i,T_{i}})_{i=1}^{N}$ using a sample that holds out final-period observations. Figure 2 compares the resulting predictions and prediction errors from the individual MLE and the EB methods. The one-period-ahead prediction of $Y_{i,T_{i}}$ is computed as
$\hat{Y}_{i,T_{i}}$ $=\hat{a}_{i}+X_{2,iT_{i}}\hat{b}_{i}+\hat{\rho}_{i}Y_{iT_{i}-1}-\widehat{\rho_{i}a_{i}}-X_{2,iT_{i}-1}\widehat{\rho_{i}b_{i}}$, (17)
where the hats indicate either EB or individual-level MLE estimates. The EB estimators are obtained by reestimating the NPMLE using the sample that excludes the last-period observations, i.e., $\{(Y_{it},X_{2,it}):i=1,\ldots,N;t=1,\ldots,T_{i}-1\}$. As shown in Figure 1, both predictions and prediction errors from EB exhibit substantial shrinkage compared to those based on MLE, resulting in variance reductions of approximately $25\%$ for predictions and $20\%$ for prediction errors.
Figure 2: Pseudo-out-of-sample predictions and prediction errors
(a) Predictions: $\hat{Y}^{\mathrm{MLE}}_{i,T_{i}}$
(b) Predictions: $\hat{Y}^{\mathrm{EB}}_{i,T_{i}}$
(c) Errors: $\hat{Y}^{\mathrm{MLE}}_{i,T_{i}}-Y_{i,T_{i}}$
(d) Errors: $\hat{Y}^{\mathrm{EB}}_{i,T_{i}}-Y_{i,T_{i}}$
Notes: The reported standard deviations of the prediction errors are computed after excluding a single outlier observation whose prediction error falls below 4.5-4.5 under both methods; if this observation is included, the standard deviations for the MLE and EB errors are $0.456$ and $0.411$, respectively.
We now discuss the estimation results concerning the joint distribution of the random components. As shown in Table 1, a notable feature is the negative correlation between $a_{i}$ and $b_{i}$. To further examine this relationship, the left panel of Figure 3 plots the EB estimate pairs $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})$ for $N=938$ individuals, revealing a clear negative association between the individual intercepts $\hat{a}_{i}^{\mathrm{EB}}$ and slopes $\hat{b}_{i}^{\mathrm{EB}}$. [^14] For completeness, Figure 4 in Appendix C presents all pairwise scatter plots of the EB estimates for $\theta_{i}$. To explore the experience profile in the cross-section, we use the estimated moments in Table 1 to compute the cross-sectional prior variance of $a_{i}+b_{i}x$, which captures the deviation of an individual’s income trajectory at $X_{2,it}=x$ from the average. The mapping $x\mapsto\mathbb{V}_{\hat{G}}(a_{i}+b_{i}x)$, shown in the right panel of Figure 3, indicates that the variance declines up to about $x=1$ (ten years of experience) and increases thereafter. This non-monotonic pattern arises from the negative covariance between $a_{i}$ and $b_{i}$. Interestingly, a similar U-shaped relationship between the log variance of residual earnings and experience is documented in Mincer (1974, Chapter 6, p.104), who attributes it to “a weak correlation between post-school investment ratios and earning capacity.”
Figure 3: EB Estimates of $(a_{i},b_{i})$ and Cross-Sectional Variance of Earnings
(a) Pairs $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})$.
(b) Variance profile by experience.
Note: The left panel plots the EB estimates $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})_{i=1}^{N}$. The right panel displays the cross-sectional variance of earnings as a function of years of experience, $\text{Exp}\mapsto\mathbb{V}_{\hat{G}}(a_{i}+b_{i}\text{Exp})$.
8 Monte Carlo Experiments
We conduct Monte Carlo experiments to evaluate the performance of the EB estimator in dynamic panel models relative to (i) the oracle decision rule and (ii) the individual MLE.
The following DGPs are considered.
  1. 1.
    The HIVD model with $\theta_{i}=(a_{i},\sigma_{i}^{2},\rho_{i})$. The individual intercepts $(a_{i})_{i=1}^{N}$ are drawn from the $\mathrm{Gamma}(1/2,\sqrt{2})$ distribution independently of $(\sigma_{i}^{2},\rho_{i})$. The joint distribution of $(\sigma_{i}^{2},\rho_{i})$ assigns probability mass $1/6$ to $(\sigma_{i}^{2},\rho_{i})=(0.1,0.8)$ and $(0.3,0.2)$, and probability mass $1/3$ to $(\sigma_{i}^{2},\rho_{i})=(0.1,0.2)$ and $(0.3,0.8)$. This design features a negative correlation of 0.33-0.33 between $\rho_{i}$ and $\sigma_{i}^{2}$. The initial $u_{i0}$ is drawn from the stationary distribution $\mathcal{N}(0,\sigma_{i}^{2}/(1-\rho_{i}^{2}))$. The sample size for estimation is $(N,T)=(1000,5)$. An additional observation $Y_{iT+1}$ is generated for each unit to compute prediction errors.
  2. 2.
    The Restricted HIVD model follows the HIVD model above but imposes $\rho_{i}\equiv 0.5$ and assumes $\sigma_{i}^{2}$ has a two-point distribution satisfying $G_{*}(\sigma_{i}^{2}=0.1)=G_{*}(\sigma_{i}^{2}=0.3)=0.5$. The initial distribution of $u_{i0}$ is $\mathcal{N}(0,\sigma_{i}^{2}/(1-0.5^{2}))$. The distribution of $a_{i}$ and the sample size is the same as in HIVD. This scenario is that a researcher does not know that $\rho_{i}$ is a constant and continue to estimate the HIVD model.
  3. 3.
    The HIDVX model generates $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ from the prior distribution estimated from the PSID sample. The deterministic covariate (interpreted as experience divided by 10) is assumed to evolve according to $X_{2,it}=X_{2,i,t-1}+0.1$, with $(N,T)=(1000,13)$, approximately as in the PSID sample. The initial value $X_{2,i1}$ is drawn from a discrete distribution with probabilities $\mathbb{P}(X_{2,i1}=0.3)=0.25$, $\mathbb{P}(X_{2,i1}=0.4)=0.05$, $\mathbb{P}(X_{2,i1}=0.5)=\mathbb{P}(X_{2,i1}=0.6)=0.1$, and $\mathbb{P}(X_{2,i1}=0.7)=0.5$.
For the first two models, the WFR parameters are $\eta=0.1$, $\overline{n}=2000$, and $\mathtt{tol}=10^{-4}$. For the more complex HIVDX, $\eta=0.005$, $\overline{n}=10000$ without applying the early stopping criterion to ensure stable and reliable convergence of the algorithm.
Let $\hat{\varepsilon}_{i}=\hat{\tau}_{i}-\tau_{i}$ denote the estimation error for unit $i$, where $\tau_{i}$ is one of the model parameters, or the one step ahead prediction. For each method and replication, we compute
$\mathrm{Bias}=\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i},\qquad\mathrm{SD}=\left(\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\varepsilon}_{i}-\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}\right)^{2}\right)^{1/2},\qquad\mathrm{RMSE}=\left(\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}^{2}\right)^{1/2}.$
Additionally, a measure of predictability, denoted by R2, is computed as one minus the ratio of the mean squared prediction error to the sample variance of $Y_{iT+1}$.
Table 2 reports results for the HIVD model and its restricted variant. In both designs, the oracle estimator achieves the lowest estimation error, but the EB estimator improves upon the individual MLE with uniformly smaller RMSE across all parameters. The gains are particularly pronounced for $\rho_{i}$. In line with this theoretical prediction, the gains from EB are even larger in the restricted HIVD design, where the prior for $\rho_{i}$ is more informative. Most of the RMSEs of the EB estimator are within 10% of those of the oracle.
Table 3 presents the results for the HIVDX model. Similarly to Table 2, the EB estimator outperforms the individual MLE across all parameters. For both $a_{i}$ and $b_{i}$, the EB method achieves MSEs that are approximately 55% of those of the MLE. The EB estimator substantially reduces the MSE for $\rho_{i}$ as well, primarily attributable to bias reduction. It is interesting to note that the Monte Carlo results for the bias of the MLE of $\rho_{i}$ echo the pattern observed in the empirical application in Section 7. In terms of out-of-sample prediction accuracy, R2 increases from 0.591 for the MLE to 0.662 for EB.
Finally, Table 4 shows the prior moment estimation results in the HIVDX design. The first moments of $\theta_{i}$ are generally well estimated, with the exception of a modest downward bias in the estimate of $\mathbb{E}_{G_{*}}[\rho_{i}]$. The variances of $\theta_{i}$ tend to be somewhat overestimated, which is expected given that estimation noise generally inflates the estimated variance of $\theta_{i}$ relative to the true variance. The negative covariance between $a_{i}$ and $b_{i}$ is estimated with a downward bias but the sign is reliably recovered.
Table 2: Errors from 500 Monte Carlo Replications
Estimator
Errors DGP Metric Oracle MLE EB
$\hat{a}_{i}-a_{i}$ HIVD Bias 0.004-0.004 $0.000$ 0.002-0.002
SD $0.378$ $0.449$ $0.414$
RMSE $0.378$ $0.449$ $0.414$
HIVDR Bias $0.001$ $0.000$ 0.003-0.003
SD $0.303$ $0.346$ $0.339$
RMSE $0.303$ $0.346$ $0.339$
$\hat{\sigma}_{i}^{2}-\sigma_{i}^{2}$ HIVD Bias $0.000$ 0.077-0.077 0.006-0.006
SD $0.078$ $0.117$ $0.085$
RMSE $0.078$ $0.141$ $0.086$
HIVDR Bias $0.000$ 0.077-0.077 0.007-0.007
SD $0.079$ $0.118$ $0.085$
RMSE $0.079$ $0.141$ $0.086$
$\hat{\rho}_{i}-\rho_{i}$ HIVD Bias 0.003-0.003 0.552-0.552 0.062-0.062
SD $0.273$ $0.482$ $0.323$
RMSE $0.273$ $0.733$ $0.330$
HIVDR Bias $0.000$ 0.546-0.546 0.048-0.048
SD $0.000$ $0.447$ $0.174$
RMSE $0.000$ $0.706$ $0.182$
$\hat{Y}_{iT+1}-Y_{iT+1}$ HIVD Bias 0.001-0.001 $0.000$ $0.000$
SD $0.485$ $0.560$ $0.490$
RMSE $0.485$ $0.560$ $0.490$
$\mathrm{R}^{2}$ $0.825$ $0.766$ $0.821$
HIVDR Bias 0.001-0.001 $0.000$ 0.001-0.001
SD $0.471$ $0.562$ $0.477$
RMSE $0.471$ $0.562$ $0.477$
$\mathrm{R}^{2}$ $0.824$ $0.749$ $0.819$
Notes: Each row reports averages computed from $500$ replications. Oracle refers to the true posterior mean, and MLE corresponds to the individual MLE.
Table 3: Errors from 500 Monte Carlo Replications (HIVDX)
Estimator
Errors Metric Oracle MLE EB
$\hat{a}_{i}-a_{i}$ Bias $0.001$ 0.001-0.001 0.001-0.001
SD $0.244$ $0.426$ $0.318$
RMSE $0.244$ $0.426$ $0.318$
$\hat{b}_{i}-b_{i}$ Bias 0.001-0.001 $0.001$ $0.001$
SD $0.188$ $0.338$ $0.251$
RMSE $0.188$ $0.338$ $0.251$
$\hat{\sigma}_{i}^{2}-\sigma_{i}^{2}$ Bias $0.000$ 0.020-0.020 0.006-0.006
SD $0.047$ $0.066$ $0.064$
RMSE $0.047$ $0.069$ $0.064$
$\hat{\rho}_{i}-\rho_{i}$ Bias 0.002-0.002 0.326-0.326 0.082-0.082
SD $0.212$ $0.308$ $0.275$
RMSE $0.212$ $0.448$ $0.288$
$\hat{Y}_{iT+1}-Y_{iT+1}$ Bias $0.000$ $0.001$ $0.001$
SD $0.299$ $0.341$ $0.314$
RMSE $0.299$ $0.341$ $0.314$
$\mathrm{R}^{2}$ $0.692$ $0.591$ $0.662$
Notes: Each row reports averages computed from $500$ replications. Oracle refers to the true posterior mean, and MLE corresponds to the individual MLE.
Table 4: Monte Carlo Results for Moment Estimates of $G_{*}$ (HIVDX)
Panel A. Means
$\mathbb{E}_{\hat{G}}[a_{i}]$ $\mathbb{E}_{\hat{G}}[b_{i}]$ $\mathbb{E}_{\hat{G}}[\sigma_{i}^{2}]$ $\mathbb{E}_{\hat{G}}[\rho_{i}]$
Truth $0.013$ 0.017-0.017 $0.079$ $0.422$
Bias 0.001-0.001 $0.001$ 0.006-0.006 0.082-0.082
SD $0.022$ $0.017$ $0.005$ $0.016$
RMSE $0.023$ $0.018$ $0.011$ $0.080$
Panel B. Variances
$\mathbb{V}_{\hat{G}}(a_{i})$ $\mathbb{V}_{\hat{G}}(b_{i})$ $\mathbb{V}_{\hat{G}}(\sigma_{i}^{2})$ $\mathbb{V}_{\hat{G}}(\rho_{i})$
Truth $0.205$ $0.101$ $0.023$ $0.105$
Bias $0.054$ $0.030$ 0.007-0.007 $0.017$
SD $0.037$ $0.025$ $0.006$ $0.009$
RMSE $0.065$ $0.039$ $0.009$ $0.020$
Panel C. Pairwise Covariances
$(a_{i},b_{i})$ $(a_{i},\sigma_{i}^{2})$ $(a_{i},\rho_{i})$ $(b_{i},\sigma_{i}^{2})$ $(b_{i},\rho_{i})$ $(\rho_{i},\sigma_{i}^{2})$
Truth 0.100-0.100 0.011-0.011 0.004-0.004 0.010-0.010 $0.007$ 0.001-0.001
Bias 0.038-0.038 0.003-0.003 $0.005$ $0.007$ 0.004-0.004 $0.003$
SD $0.028$ $0.011$ $0.009$ $0.010$ $0.006$ $0.002$
RMSE $0.047$ $0.010$ $0.011$ $0.012$ $0.008$ $0.003$
Notes: The row labeled ‘Truth’ reports the true values of the corresponding moments. Results are based on 500 replications.
9 Conclusions
We have proposed a nonparametric empirical Bayes framework for analyzing short-panel data with rich forms of unobserved heterogeneity. The analysis is a significant generalization of classical G-modeling to accommodate heterogeneous slopes and non-spherical error structures. Identification, consistency, and optimality results are established under general conditions, with primitive sufficient conditions derived for specific cases of interest. However, we have not developed any accompanying inference methods. Recent contributions such as Armstrong, Kolesár, and Plagborg-Møller (2022) and Ignatiadis and Wager (2022) can be useful for future research.
Appendix A Proofs
A.1 Proofs of Main Results
Proof of Theorem 1.
We divide the proof into two parts. Part (i) establishes identification of the marginal distribution of $\delta_{i}$; part (ii) builds on this to show that the full mixing distribution $G_{*}$ is identified.
Part (i). Take any solution $G_{o}\in\mathcal{G}$ to (7), which must satisfy $f_{G_{o}}(Y_{i},X_{i})=f_{G_{*}}(Y_{i},X_{i})$ almost surely. Let $H_{*}$ and $H_{o}$ denote the marginal distributions of $\delta_{i}$ induced by $G_{*}$ and $G_{o}$, respectively. In what follows, we show that $H_{*}=H_{o}$.
Let $\boldsymbol{X}$} and $\boldsymbol{M}$ be given as in Assumption 2. Conditioning on $X_{i}={\boldsymbol{X}}$ and premultiplying both sides of (2) by $\boldsymbol{M}^{\prime}$, we obtain $\boldsymbol{M}^{\prime}Y_{i}=\boldsymbol{M}^{\prime}P_{i}e_{i}$. Taking conditional characteristic functions of $\boldsymbol{M}^{\prime}Y_{i}$ according to $f_{G_{o}}(\cdot\,|\,{\boldsymbol{X}})$ and $f_{G_{*}}(\cdot\,|\,{\boldsymbol{X}})$, respectively, we have
$\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}\boldsymbol{M}^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]$ $=\int e^{-t^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}t/2}dH_{o}(\delta),$
$\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}\boldsymbol{M}^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]$ $=\int e^{-t^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}t/2}dH_{*}(\delta).$
Equating $\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}\boldsymbol{M}^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]$ with $\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}\boldsymbol{M}^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]$, we have
$\int_{\mathcal{K}_{\delta}}\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{1/2}}{(2\pi)^{(T-{d_{\beta}})/2}}e^{-t^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}t/2}\cdot|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}dH_{o}(\delta)$
$=$ $\ \int_{\mathcal{K}_{\delta}}\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{1/2}}{(2\pi)^{(T-{d_{\beta}})/2}}e^{-t^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}t/2}\cdot|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}dH_{*}(\delta),$ (18)
which can be rewritten as
$\mathcal{T}[0,(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\tilde{H}_{o}](t)=\mathcal{T}[0,(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\tilde{H}_{*}](t),\quad\forall t\in\mathbb{R}^{T-{d_{\beta}}},$
with $\mathcal{T}$ defined before Lemma 1. The probability measures $\tilde{H}_{o}$ and $\tilde{H}_{*}$ are defined via
$d\tilde{H}_{o}(\delta):=\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}}{C}dH_{o}(\delta),\quad d\tilde{H}_{*}(\delta):=\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}}{C}dH_{*}(\delta),$
where the norming constant $C>0$ is
$C=\int_{\mathcal{K}_{\delta}}|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}dH_{*}(\delta)=\int_{\mathcal{K}_{\delta}}|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}dH_{*}(\delta).$
This equality follows from integrating both sides of (A.1) with respect to $t\in\mathbb{R}^{T-{d_{\beta}}}$, ensuring that both $\tilde{H}_{*}$ and $\tilde{H}_{o}$ are probability measures. We are prepared to apply Lemma 1. Since $\delta\mapsto(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1}$ is injective by Assumption 2(ii), (30) is satisfied by $\Lambda(\mathcal{K}_{\delta})$ which consists only of $(\lambda_{1}(\delta),\lambda_{2}(\delta)):=(0,(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1})$. By Lemma 1, it follows that $\tilde{H}_{*}=\tilde{H}_{o}$, and hence $H_{*}=H_{o}$.
Part (ii). Next, we show that the conditional distribution of ${\boldsymbol{X}}\beta_{i}$ given $\delta_{i}=\delta$ is identified for each $\delta$ in the support of $H_{*}$.
Let $\boldsymbol{X}$} and $\boldsymbol{M}$ be again given as in Assumption 2. Computing the conditional characteristic functions of $Y_{i}$ given $X_{i}={\boldsymbol{X}}$ according to the densities $f_{G_{o}}(\cdot\,|\,{\boldsymbol{X}})$ and $f_{G_{*}}(\cdot\,|\,{\boldsymbol{X}})$, respectively, we obtain:
$\varphi_{o}(t)$ $:=\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]=\int\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]e^{-t^{\prime}P(\delta)P(\delta)^{\prime}t/2}dH_{o}(\delta),$
$\varphi_{*}(t)$ $:=\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}Y_{i}}|X_{i}={\boldsymbol{X}}]=\int\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]e^{-t^{\prime}P(\delta)P(\delta)^{\prime}t/2}dH_{*}(\delta).$
These function coincide with each other since $f_{G_{o}}=f_{G_{*}}$ By Part (i), we also have $H_{o}=H_{*}$. Fix $t\in\mathbb{R}^{T}$, and let
$\psi(\delta):=\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]-\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta].$
Equating $\varphi_{o}(t+\boldsymbol{M}s)$ with $\varphi_{*}(t+\boldsymbol{M}s)$, we then obtain
$\int\psi(\delta)\exp\left(-\frac{1}{2}\|P(\delta)^{\prime}(\boldsymbol{M}s+t)\|^{2}\right)dH_{*}(\delta)=0$ (19)
for all $s\in\mathbb{R}^{T-{d_{\beta}}}$.
For each $\delta\in\mathcal{K}_{\delta}$, we write
$t=\boldsymbol{M}\alpha_{\delta}+e_{\delta},$
where $\alpha_{\delta}=(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}t$ is chosen so that $\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}e_{\delta}=0$. This allows us to rewrite (19) as
$0$ $=\int\psi(\delta)\exp\left(-\frac{1}{2}\|P(\delta)^{\prime}e_{\delta}\|^{2}-\frac{1}{2}(s+\alpha_{\delta})^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}(s+\alpha_{\delta})\right)dH_{*}(\delta)$
$={\int}\begin{aligned} &(2\pi)^{(T-{d_{\beta}})/2}|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}\psi(\delta)\exp\left(-\frac{1}{2}\|P(\delta)^{\prime}e_{\delta}\|^{2}\right)\\ &\quad\times\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{1/2}}{(2\pi)^{(T-{d_{\beta}})/2}}\exp\left(-\frac{1}{2}(s+\alpha_{\delta})^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}(s+\alpha_{\delta})\right)dH_{*}(\delta)\end{aligned}$ (20)
for all $s\in\mathbb{R}^{r}$. Define a complex Borel measure
$d\mu(\delta):=(2\pi)^{(T-{d_{\beta}})/2}|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{-1/2}\psi(\delta)\exp\left(-\frac{1}{2}\|P(\delta)^{\prime}e_{\delta}\|^{2}\right)dH_{*}(\delta).$
Our next goal is to establish that the real and imaginary parts of $\mu$ are both zero. By construction of $\mu$, (A.1) implies that
$\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu](s)$
$=$ $\ \int\frac{|\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}|^{1/2}}{(2\pi)^{(T-{d_{\beta}})/2}}\exp\left(-\frac{1}{2}(s+\alpha_{\delta})^{\prime}\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}(s+\alpha_{\delta})\right)d\mu(\delta)=0$ (21)
for all $s\in\mathbb{R}^{T-{d_{\beta}}}$. Using the Jordan decomposition theorem, we decompose $\mu$ into
$\mu=\mu_{+}^{R}-\mu_{-}^{R}+\mathrm{i}(\mu_{+}^{I}-\mu_{-}^{I}),$
where $\mu_{\pm}^{R}$ are mutually singular non-negative Borel measures, representing the positive and negative real parts of $\mu$, respectively. The same holds for the positive and negative imaginary parts $\mu_{\pm}^{I}$. By linearity of $\mathcal{T}$ with respect to $\mu$, it follows from (A.1) that
$\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu^{R}_{+}](s)$ $=\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu^{R}_{-}](s),$ (22)
$\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu_{I}^{+}](s)$ $=\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu_{I}^{-}](s).$ (23)
Integrating both sides of (22) with respect to $s\in\mathbb{R}^{T-{d_{\beta}}}$, by Fubini’s theorem, it follows
$\mu^{R}_{+}(\mathcal{K}_{\delta})$ $=\int_{\mathbb{R}^{T-{d_{\beta}}}}\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu^{R}_{+}](s)ds$
$=\int_{\mathbb{R}^{T-{d_{\beta}}}}\mathcal{T}[-\alpha_{\delta},(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1},\mu^{R}_{-}](s)ds=\mu^{R}_{-}(\mathcal{K}_{\delta}).$
We examine two possible cases. If $\mu_{-}^{R}(\mathcal{K}_{\delta})=\mu_{+}^{R}(\mathcal{K}_{\delta})=0$, the real part of $\mu$ vanishes trivially. Otherwise, we may without loss of generality assume that both $\mu_{-}^{R}$ and $\mu_{+}^{R}$ are probability measures. Lemma 1 then implies that $\mu^{R}_{+}=\mu^{R}_{-}$. However, since $\mu^{R}_{+}$ and $\mu^{R}_{-}$ are mutually singular, this is only possible when both $\mu^{R}_{+}$ and $\mu^{R}_{-}$ have total mass of zero, which leads to a contradiction. Thus, we conclude $\mu_{-}^{R}=\mu_{+}^{R}\equiv 0$, and hence the real part of $\mu$ is identically zero. The same applies to $\mu_{I}^{+}$ and $\mu_{I}^{-}$. It follows $\mu\equiv 0$, thus implying that $\psi(\delta)=0$ $H_{*}$-a.e. on $\mathcal{K}_{\delta}$.
Thus far, we have established that
$\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]-\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]=0$
for $H_{*}$-a.e. $\delta$ for each fixed $t\in\mathbb{R}^{T}$. Define the function
$\omega(t,\delta):=\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]-\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]$
which admits a jointly measurable version, and hence can be regarded as Borel-measurable in $(t,\delta)$. By Fubini’s theorem, we have
$0=\int_{\mathbb{R}^{T}}H_{*}\left(\{\delta\in\mathcal{K}_{\delta}:\omega(t,\delta)\neq 0\}\right)dt$ $=(\operatorname{Leb}\times H_{*})\left(\{(t,\delta)\in\mathbb{R}^{T}\times\mathcal{K}_{\delta}:\omega(t,\delta)\neq 0\}\right)$
$=\int_{\mathcal{K}_{\delta}}\operatorname{Leb}\left(\{t\in\mathbb{R}^{T}:\omega(t,\delta)\neq 0\}\right)dH_{*}(\delta),$
where $\operatorname{Leb}\times H_{*}$ denotes the product measure of the Lebesgue measure and $H_{*}$. It follows that $\operatorname{Leb}(\{t:\omega(t,\delta)\neq 0\})=0$ for $H_{*}$-a.e. $\delta\in\mathcal{K}_{\delta}$. Observe that $\omega(\cdot,\delta)$ is a uniformly continuous function of $t\in\mathbb{R}^{T}$ for $\delta\in\mathcal{K}_{\delta}$. Since the set of all zeros of $\omega(\cdot,\delta)$ is dense in $\mathbb{R}^{T}$, $\omega(\cdot,\delta)$ must vanish everywhere. It follows that for $H_{*}$-a.e. $\delta$,
$\mathbb{E}_{G_{o}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta]=\mathbb{E}_{G_{*}}[e^{\mathrm{i}t^{\prime}{\boldsymbol{X}}\beta_{i}}|\delta_{i}=\delta],\quad\forall t\in\mathbb{R}^{T}.$
This establishes that the conditional distribution of ${\boldsymbol{X}}\beta_{i}$ given $\delta_{i}$ is identified. Finally, since $\boldsymbol{X}$} has full column rank, this is equivalent to identifying the conditional distribution of $\beta_{i}$ given $\delta_{i}$, implying that $G_{*}$ is identified from the conditional distribution $\beta_{i}\,|\,\delta_{i}$ and the marginal distribution of $\delta_{i}$. ∎
Proof of Theorem 2.
We verify a multivariate version of Assumptions 1-5 in Kiefer and Wolfowitz (1956).
KW Assumption 1 is trivially met, since $Y\mapsto f_{G}(Y,X)$ is a legitimate density with respect to the Lebesgue measure on $\mathbb{R}^{T}$. KW Assumption 2 requires
$f_{\tilde{\mu}}(Y,X)\to f_{\mu}(Y,X)\quad\text{as}\quad{d}(\tilde{\mu},\mu)\to 0$
for any $\mu$ and $\mathbb{P}$-a.e. $(Y,X)$, which is established in Lemma 2. KW Assumptions 3 and 4 are verified in Lemma 3(ii) and (iii), respectively. Finally, KW Assumption 5 follows from the fact that $\mathbb{E}[|\log f_{G_{*}}(Y_{i},X_{i})|]<\infty$ and the random variable $\log\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i})$ is uniformly bounded above for all $\mu\in\bar{\mathcal{G}}$ and $\varepsilon>0$, as established in Lemma 3(i) and (ii). ∎
Proof of Theorem 3.
We first show that $\mathbb{E}[N^{-1}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{p}]\to 0$ as $N\to\infty$. Let $\epsilon>0$ be fixed, and let $M>0$ be given as in Assumption 4(ii). Consider the following decomposition
$\tau_{i}=\tau_{M,i}+r_{M,i},$
where
$\tau_{M,i}\equiv\tau_{M}(\theta_{i}):=\frac{M}{\max\{|\tau(\theta_{i})|,M\}}\tau(\theta_{i})\equiv\begin{cases}\tau(\theta_{i})&\text{if }|\tau(\theta_{i})|\leq M\\ \frac{M}{|\tau(\theta_{i})|}\tau(\theta_{i})&\text{if }|\tau(\theta_{i})|>M\end{cases}$
denotes the truncation of $\tau_{i}$ at level $M$, and $r_{M,i}=r_{M}(\theta_{i}):=\tau(\theta_{i})-\tau_{M}(\theta_{i})$ is the remainder term. By construction, part (i) of Assumption 4 implies that $\tau_{M}(\theta)$ is a bounded continuous function of $\theta$, and part (ii) implies
$\limsup_{N\to\infty}\mathbb{E}\int_{\Theta}\|r_{M}(\theta)\|^{p}d\hat{G}_{N}(\theta)\leq\epsilon.$
Write $\tau_{M,i}^{*}$ and $\hat{\tau}^{\mathrm{EB}}_{M,i}$ for the oracle and the EB decision rules for $\tau_{M,i}$, and similarly for $r_{M,i}^{*}$ and $\hat{r}_{M,i}$. Using the fact that ${\tau}^{*}_{i}=\tau_{M,i}^{*}+r_{M,i}^{*}$ and $\hat{\tau}^{\mathrm{EB}}_{i}=\hat{\tau}^{\mathrm{EB}}_{M,i}+\hat{r}^{{\mathrm{EB}}}_{M,i}$, we have
$\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{p}$ $\leq C_{p}\left(\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{M,i}-\tau_{M,i}^{*}|^{p}+\frac{1}{N}\sum_{i=1}^{N}|\hat{r}_{M,i}-r_{M,i}^{*}|^{p}\right)$
$=:C_{p}(I_{1}+I_{2})$
for some $C_{p}<\infty$ that only depends on $p$. In the following steps, we show that both $\mathbb{E}[I_{1}]$ and $\mathbb{E}[I_{2}]$ converge to zero.
Let $M_{0}>0$ be a sufficiently large number such that $\mathbb{P}(\|Y_{i}\|\leq M_{0},\,M_{0}^{-1}{I}_{{d_{\beta}}}\leq X_{i}^{\prime}X_{i}\leq M_{0}{I}_{{d_{\beta}}})\geq 1-\epsilon/(2M)^{p}$. Let $\mathcal{R}=\{(Y,X):\|Y\|\leq M_{0},\,M_{0}^{-1}{I}_{{d_{\beta}}}\leq X^{\prime}X\leq M_{0}{I}_{{d_{\beta}}}\}$, and let $\mathcal{R}_{i}=1\{\|Y_{i}\|\leq M_{0},\,M_{0}^{-1}{I}_{{d_{\beta}}}\leq X_{i}^{\prime}X_{i}\leq M_{0}{I}_{{d_{\beta}}}\}$ be the corresponding indicator for unit $i$. We decompose $I_{1}$ into two terms:
$I_{11}=\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{M,i}-\tau_{M,i}^{*}|^{p}(1-\mathcal{R}_{i}),\quad I_{12}=\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{M,i}-\tau_{M,i}^{*}|^{p}\mathcal{R}_{i}.$
It is straightforward to see that
$\mathbb{E}[I_{11}]\leq(2M)^{p}\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(1-\mathcal{R}_{i})\right]\leq\epsilon$
by construction of $\mathcal{R}_{i}$. On the other hand,
$\mathbb{E}[I_{12}]\leq\mathbb{E}\sup_{i:(Y_{i},X_{i})\in\mathcal{R}}|\hat{\tau}^{\mathrm{EB}}_{M,i}-\tau_{M,i}^{*}|^{p}\leq\mathbb{E}\sup_{(Y,X)\in\mathcal{R}}|\hat{\tau}^{\mathrm{EB}}_{M}(Y,X)-\tau_{M}^{*}(Y,X)|^{p}\to 0$
by Lemma 4 and the bounded convergence theorem (Royden and Fitzpatrick, 2010). These results imply that $\limsup_{N\to\infty}\mathbb{E}[I_{1}]\leq\epsilon$.
To show $\limsup_{N\to\infty}\mathbb{E}[I_{2}]\leq\epsilon$, first observe that
$I_{2}$ $\leq C_{p}\left(\frac{1}{N}\sum_{i=1}^{N}|\hat{r}^{{\mathrm{EB}}}_{M,i}|^{p}+\frac{1}{N}\sum_{i=1}^{N}|r_{M,i}^{*}|^{p}\right)=:C_{p}(I_{21}+I_{22}).$
For the first term, we have, by the Jensen’s inequality,
$|\hat{r}^{{\mathrm{EB}}}_{M,i}|^{p}\leq\mathbb{E}_{\hat{G}}[|r_{M}(\theta_{i})|^{p}\,|\,Y_{i},X_{i}]=\frac{\int_{\Theta}|r_{M}(\theta)|^{p}\ell(Y_{i}\,|\,X_{i},\theta)d\hat{G}_{N}(\theta)}{f_{\hat{G}}(Y_{i},X_{i})},\quad i=1,\ldots,N.$
This implies
$\frac{1}{N}\sum_{i=1}^{N}|\hat{r}^{{\mathrm{EB}}}_{M,i}|^{p}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\frac{\int_{\Theta}|r_{M}(\theta)|^{p}\ell(Y_{i}\,|\,X_{i},\theta)d\hat{G}_{N}(\theta)}{f_{\hat{G}}(Y_{i},X_{i})}$
$=\int_{\Theta}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta).$
The first-order conditions for the NPMLE state that $N^{-1}\sum_{i=1}^{N}{\ell(Y_{i}\,|\,X_{i},\theta)}/{f_{\hat{G}}(Y_{i},X_{i})}=1$ for $\hat{G}_{N}$-a.e. $\theta$, which imply
$\int_{\Theta}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta)=\int_{\Theta}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta).$
It follows that $\limsup_{N\to\infty}\mathbb{E}[I_{21}]\leq\epsilon$ by Assumption 4(ii).
Turning to the term $I_{22}$, we have, by Jensen’s inequality,
$I_{22}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{G_{*}}[|r_{M}(\theta_{i})|^{p}\,|\,Y_{i},X_{i}],$
which implies that $\limsup_{N\to\infty}\mathbb{E}[I_{22}]\leq\mathbb{E}_{G_{*}}[|r_{M}(\theta_{i})|^{p}]$. Since $\theta\mapsto|r_{M}(\theta)|^{p}$ is a nonnegative, continuous function of $\theta$ and $\hat{G}_{N}$ converges weakly to $G_{*}$ almost surely, Portmanteau theorem implies that
$\int_{\Theta}|r_{M}(\theta)|^{p}dG_{*}(\theta)\leq\liminf_{N\to\infty}\int_{\Theta}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta)\quad\text{a.s.}$
By Fatou’s lemma, $\mathbb{E}[\liminf_{N\to\infty}\int_{\Theta}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta)]\leq\liminf_{N\to\infty}\mathbb{E}[\int_{\Theta}|r_{M}(\theta)|^{p}d\hat{G}_{N}(\theta)]$. By Assumption 4(ii), this implies that $\limsup_{N\to\infty}\mathbb{E}[I_{22}]\leq\epsilon$.
Putting these pieces together, we conclude that
$\limsup_{N\to\infty}\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{p}\right]\leq 3C_{p}\epsilon.$
Since $\epsilon>0$ was arbitrarily chosen, we have $\mathbb{E}[N^{-1}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{p}]\to 0$.
To prove the regret consistency of $\hat{\tau}^{\mathrm{EB}}_{i}$, observe that Assumption 4 holding for some $p\in[2,\infty)$ automatically implies that it also holds for $p=2$. By the same argument, we have $\mathbb{E}[N^{-1}\sum_{i=1}^{N}|\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i}|^{2}]\to 0$, from which regret consistency follows. ∎
Proof of Proposition 1.
It suffices to verify Assumption 2 and it is verified in the main text. See Section 3.1.1 for details. ∎
Proof of Proposition 2.
Again, it suffices to verify Assumption 2. See Appendix B.1 for details. ∎
Proof of Proposition 3.
By assumption, there exists $\boldsymbol{x}=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{x}_{3},\boldsymbol{x}_{4})\in\operatorname{supp}(X_{2,i})$ such that either $\Delta\boldsymbol{x}_{2}=\pm\Delta\boldsymbol{x}_{3}\neq 0$ or $\Delta\boldsymbol{x}_{2}\neq\Delta\boldsymbol{x}_{4}$ where $\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t}-\boldsymbol{x}_{t-1}$. Upon inspecting the proof of Proposition 4, we find that the assumption $T\geq 5$ is explicitly invoked only in Cases 1(c), 2(c), and 3(c); in the other cases, $T=4$ suffices for identification. Further inspection reveals that Case 1(c) requires $(\Delta\boldsymbol{x}_{2},\Delta\boldsymbol{x}_{3},\Delta\boldsymbol{x}_{4})=(0,z,0)$ for some $z\neq 0$. Similarly, Cases 2(c) and 3(c) require that $(\Delta\boldsymbol{x}_{2},\Delta\boldsymbol{x}_{3},\Delta\boldsymbol{x}_{4})$ take the forms $(z,0,z)$ for some $z\neq 0$, and $(z,z^{\prime},z)$ for some $\neq\pm z^{\prime}, z,z$$z^{\prime}\neq 0,$ respectively. Since the given $\boldsymbol{x}$ belongs to none of these cases, the identification follows from the proof of Proposition 4. ∎
Proof of Proposition 4.
We prove Proposition 4 by verifying Assumption 2. Let $\Delta a_{t}:=a_{t+1}-a_{t}$ denote the forward first difference of $a_{t}$. By the assumption that $\Delta X_{2,it}$ is not identically zero, there is $\boldsymbol{x}=(\boldsymbol{x}_{t})_{t=1}^{T}\in\operatorname{supp}(X_{2,i})$ such that $\Delta{\boldsymbol{x}}=(\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t})_{t=1}^{T-1}\neq 0$. For notational ease, write $\boldsymbol{z}=\Delta{\boldsymbol{x}}\in\mathbb{R}^{T-1}$. Conditioning on $X_{2,i}=\boldsymbol{x}$, we obtain, after first-differencing,
$\Delta Y_{it}=b_{i}\boldsymbol{z}_{t}+\Delta u_{it},$
which yields
$\frac{\Delta Y_{it}}{\boldsymbol{z}_{t}}=b_{i}+\frac{\Delta u_{it}}{\boldsymbol{z}_{t}}\quad\text{if}\quad\boldsymbol{z}_{t}\neq 0,$
and
$\Delta Y_{it}=\Delta u_{it}\quad\text{if}\quad\boldsymbol{z}_{t}=0.$
Let $\mathcal{T}:=\{t:\boldsymbol{z}_{t}\neq 0\}$ be the set of indices at which $\boldsymbol{z}_{t}\neq 0$, with $t(1)<\cdots<t(K)$ being the increasing enumeration of $\mathcal{T}$ and $K=|\mathcal{T}|$. The annihilator matrix $\boldsymbol{M}$ can be chosen so that
$(\boldsymbol{M}^{\prime}Y_{i})_{k}=\frac{\Delta u_{i{t(k+1)}}}{\boldsymbol{z}_{t(k+1)}}-\frac{\Delta u_{i{t(k)}}}{\boldsymbol{z}_{t(k)}}\quad\text{for}\ k=1,\ldots,K-1,$
for the first K1K-1 elements, and
$((\boldsymbol{M}^{\prime}Y_{i})_{K},\ldots,(\boldsymbol{M}^{\prime}Y_{i})_{T-1})=(\Delta u_{it})_{t\in\mathcal{T}^{c}}.$
Our objective is to determine $\delta=(\sigma^{2},\rho)$ from the covariance $\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}=\mathbb{V}(\boldsymbol{M}^{\prime}Y_{i}\,|\,\delta_{i}=\delta)$ of the normalized second differences $\frac{\Delta u_{i{t(k+1)}}}{\boldsymbol{z}_{t(k+1)}}-\frac{\Delta u_{i{t(k)}}}{\boldsymbol{z}_{t(k)}}$ for $k=1,\ldots,K-1$ and the first differences $\Delta u_{it}$ for $t\in\mathcal{T}^{c}$.
The argument proceeds by nested case analysis. We first consider Case 1.
Case 1: (Two or more zeros in $\boldsymbol{z}$.) Assume $|\mathcal{T}^{c}|\geq 2$. For any $s>t$ satisfying $\boldsymbol{z}_{t}=0$ and $\boldsymbol{z}_{s}=0$, we find that
$\mathbb{V}\left(\Delta u_{it}\,|\,\delta_{i}=\delta\right)=\frac{2\sigma^{2}}{1+\rho}$
and
$\operatorname{Cov}\left(\Delta u_{it},\Delta u_{is}\,|\,\delta_{i}=\delta\right)=-\frac{\rho^{s-t-1}(1-\rho)\sigma^{2}}{1+\rho}.$
Using these equations, we can pin down the values of
$\frac{\sigma^{2}}{1+\rho}\quad\text{and}\quad\rho^{|s-t|-1}(1-\rho)\quad\forall t,s\in\mathcal{T}^{c}.$ (24)
We further divide into three subcases and show that $\delta$ can be uniquely determined in each subcase.
Case 1(a): When $s$ and $t$ are consecutive, i.e., $s=t+1$, it follows from (24) that both $1-\rho$ and $\sigma^{2}/(1+\rho)$ are uniquely determined. These together identify $\delta=(\sigma^{2},\rho)$.
Case 1(b): Suppose there are at least three elements $t_{1}<t_{2}<t_{3}$ in $\mathcal{T}^{c}$ with $d_{1}:=t_{2}-t_{1}\geq 1$ and $d_{2}:=t_{3}-t_{2}\geq 1$. Then, the values of $\rho^{d_{1}-1}(1-\rho)$, $\rho^{d_{2}-1}(1-\rho)$, and $\rho^{d_{1}+d_{2}-1}(1-\rho)$ are identified from (24). If all three are zero, this implies $\rho=0$, and vice versa. Thus, $\rho$ is identified in this case, and $\sigma^{2}$ is also identified from $\sigma^{2}/(1+\rho)$. Thus, $\delta$ is identified.
Otherwise, all three have nonzero values, and one can pin down
$\frac{\rho^{d_{1}-1}(1-\rho)\cdot\rho^{d_{2}-1}(1-\rho)}{\rho^{d_{1}+d_{2}-1}(1-\rho)}=\rho^{-1}(1-\rho)=\frac{1}{\rho}-1,$
implying that $\delta=(\sigma^{2},\rho)$ is identified in this case as well.
Case 1(c): Suppose that $|\mathcal{T}^{c}|=2$ and the elements $t,s\in\mathcal{T}^{c}$ ($t<s$) are not consecutive. Note that $|\mathcal{T}|=T-1-|\mathcal{T}^{c}|=T-3\geq 2$, and hence there are at least two time indices for which $\boldsymbol{z}_{t}\neq 0$. It must therefore be either $t-1,t+1\in\mathcal{T}$ or $s-1,s+1\in\mathcal{T}$. Without loss of generality, assume that $t-1,t+1\in\mathcal{T}$. We then observe
$\operatorname{Cov}\left(\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t-1}}}{\boldsymbol{z}_{t-1}},\Delta u_{it}\,|\,\delta_{i}=\delta\right)=-\frac{(1-\rho)\sigma^{2}}{1+\rho}\left(\frac{1}{\boldsymbol{z}_{t+1}}-\frac{1}{\boldsymbol{z}_{t-1}}\right),$
and
$\operatorname{Cov}\left(\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t-1}}}{\boldsymbol{z}_{t-1}},\Delta u_{is}\,|\,\delta_{i}=\delta\right)=-\frac{(1-\rho)\rho^{s-t-2}\sigma^{2}}{1+\rho}\left(\frac{1}{\boldsymbol{z}_{t+1}}-\frac{\rho^{2}}{\boldsymbol{z}_{t-1}}\right).$
If $\boldsymbol{z}_{t+1}\neq\boldsymbol{z}_{t-1}$, the first equation identifies ${(1-\rho)\sigma^{2}}/(1+\rho)$. Since $\sigma^{2}/(1+\rho)$ is already identified by (24), these together uniquely determine $\rho$ and $\sigma^{2}$, and hence $\delta$. If $\boldsymbol{z}_{t+1}=\boldsymbol{z}_{t-1}\neq 0$, we use the second equation to identify
$\frac{(1-\rho)\rho^{s-t-2}\sigma^{2}}{1+\rho}(1-\rho^{2})=\frac{\sigma^{2}}{1+\rho}\rho^{s-t-1}(1-\rho)\frac{1-\rho^{2}}{\rho}.$
Together with (24), this further identifies
$\frac{\rho}{1-\rho^{2}},$
which is a strictly increasing function of $\rho\in(-1,1)$. Hence, $\rho$ is identified, and so is $\delta$.
Case 2: (Only one zero in $\boldsymbol{z}$.) Let $t$ be the only index for which $\boldsymbol{z}_{t}=0$. As noted earlier in Case 1, we can then identify
$\mathbb{V}\left(\Delta u_{it}\,|\,\delta_{i}=\delta\right)=\frac{2\sigma^{2}}{1+\rho}.$
Again consider three subcases as follows.
Case 2(a): Assume that $t=1$ or t=T1t=T-1, i.e., that the only zero of $\boldsymbol{z}$ occurs either the first or last time period. Without loss of generality, let $t=1$. Since $2,3\in\mathcal{T}$, we observe that
$\mathbb{V}\left(\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}-\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}\,|\,\delta_{i}=\delta\right)=\frac{\sigma^{2}}{1+\rho}\left(\frac{2}{\boldsymbol{z}_{3}^{2}}+\frac{2}{\boldsymbol{z}_{2}^{2}}+\frac{2(1-\rho)}{\boldsymbol{z}_{3}\boldsymbol{z}_{2}}\right),$
from which we can identify
$\left(\frac{2}{\boldsymbol{z}_{3}^{2}}+\frac{2}{\boldsymbol{z}_{2}^{2}}+\frac{2(1-\rho)}{\boldsymbol{z}_{3}\boldsymbol{z}_{2}}\right).$
Clearly, this uniquely determines $\rho$, and $\sigma^{2}$ subsequently follows from $\sigma^{2}/(1+\rho)$.
Case 2(b): Assume that 1<t<T11<t<T-1 and that $\boldsymbol{z}_{t+1}\neq\boldsymbol{z}_{t-1}$, which must be nonzero. Observe that
$\operatorname{Cov}\left(\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t-1}}}{\boldsymbol{z}_{t-1}},\Delta u_{it}\,|\,\delta_{i}=\delta\right)=-\frac{(1-\rho)\sigma^{2}}{1+\rho}\left(\frac{1}{\boldsymbol{z}_{t+1}}-\frac{1}{\boldsymbol{z}_{t-1}}\right).$
From this equation, we identify $(1-\rho)\sigma^{2}/(1+\rho)$. Combined with $\sigma^{2}/(1+\rho)$, we identify $\delta$.
Case 2(c): Assume that 1<t<T11<t<T-1 and $\boldsymbol{z}_{t+1}=\boldsymbol{z}_{t-1}\neq 0$. Since $t$ is the only zero, either $\boldsymbol{z}_{t+2}\neq 0$ or $\boldsymbol{z}_{t-2}\neq 0$ must hold, depending on whether $t\geq 2$ or $t\leq T-2$. (For this, we invoke $T\geq 5$.) Without loss of generality, assume the first. Let $\boldsymbol{z}\neq 0$ denote the value shared by $\boldsymbol{z}_{t+1}=\boldsymbol{z}_{t-1}$. We use
$\mathbb{V}\left(\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t-1}}}{\boldsymbol{z}_{t-1}}\,|\,\delta_{i}=\delta\right)=\frac{2\sigma^{2}}{\boldsymbol{z}^{2}(1+\rho)}\left(2+\rho(1-\rho)\right),$
along with $\sigma^{2}/(1+\rho)$, to identify $2+\rho(1-\rho)$. Next, we employ
$\operatorname{Cov}\left(\frac{\Delta u_{i{t+2}}}{\boldsymbol{z}_{t+2}}-\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}},\Delta u_{it}\,|\,\delta_{i}=\delta\right)=-\frac{(1-\rho)\sigma^{2}}{1+\rho}\left(\frac{\rho}{\boldsymbol{z}_{t+2}}-\frac{1}{\boldsymbol{z}_{t+1}}\right),$
together with $\sigma^{2}/(1+\rho)$, to identify $\frac{(1-\rho)\rho}{\boldsymbol{z}_{t+2}}-\frac{(1-\rho)}{\boldsymbol{z}_{t+1}}$. Lastly, $\frac{(1-\rho)\rho}{\boldsymbol{z}_{t+2}}-\frac{(1-\rho)}{\boldsymbol{z}_{t+1}}$, combined with $2+\rho(1-\rho)$ identified earlier, can be used to identify $\rho$. $\sigma^{2}$ is subsequently identified from $\sigma^{2}/(1+\rho)$.
Case 3: (No zeros in $\boldsymbol{z}$.) We show that $\delta$ can be identified using the covariance structure of
$\left(\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}-\frac{\Delta u_{i{1}}}{\boldsymbol{z}_{1}},\quad\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}-\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}},\quad\frac{\Delta u_{i{4}}}{\boldsymbol{z}_{4}}-\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}\right).$
Although the last term requires $T\geq 5$, Cases 3(a) and(b) remain applicable under $T=4$, in which the first two terms are sufficient for identification.
We begin by noting that
$\mathbb{V}\left(\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t}}}{\boldsymbol{z}_{t}}\,|\,\delta_{i}=\delta\right)=\frac{2\sigma^{2}}{1+\rho}\left(1+\frac{\boldsymbol{z}_{t+1}}{\boldsymbol{z}_{t}}+\frac{\boldsymbol{z}_{t}}{\boldsymbol{z}_{t+1}}-\rho\right)\frac{1}{\boldsymbol{z}_{t+1}\boldsymbol{z}_{t}},\quad t=1,2,3.$
Consider the following three subcases.
Case 3(a): Assume that there exist $s<t$ such that $\frac{\boldsymbol{z}_{s+1}}{\boldsymbol{z}_{s}}+\frac{\boldsymbol{z}_{s}}{\boldsymbol{z}_{s+1}}\neq\frac{\boldsymbol{z}_{t+1}}{\boldsymbol{z}_{t}}+\frac{\boldsymbol{z}_{t}}{\boldsymbol{z}_{t+1}}$. We may assume $s=1$ and $t=2$. Then, we can use the equation
$\frac{\boldsymbol{z}_{2}\boldsymbol{z}_{1}\mathbb{V}\left(\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}-\frac{\Delta u_{i{1}}}{\boldsymbol{z}_{1}}\,|\,\delta_{i}=\delta\right)}{\boldsymbol{z}_{3}\boldsymbol{z}_{2}\mathbb{V}\left(\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}-\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}\,|\,\delta_{i}=\delta\right)}=\frac{1+\frac{\boldsymbol{z}_{2}}{\boldsymbol{z}_{1}}+\frac{\boldsymbol{z}_{1}}{\boldsymbol{z}_{2}}-\rho}{1+\frac{\boldsymbol{z}_{3}}{\boldsymbol{z}_{2}}+\frac{\boldsymbol{z}_{2}}{\boldsymbol{z}_{3}}-\rho}=1+\frac{(\frac{\boldsymbol{z}_{2}}{\boldsymbol{z}_{1}}+\frac{\boldsymbol{z}_{1}}{\boldsymbol{z}_{2}})-(\frac{\boldsymbol{z}_{3}}{\boldsymbol{z}_{2}}+\frac{\boldsymbol{z}_{2}}{\boldsymbol{z}_{3}})}{1+\frac{\boldsymbol{z}_{3}}{\boldsymbol{z}_{2}}+\frac{\boldsymbol{z}_{2}}{\boldsymbol{z}_{3}}-\rho}$
to pin down $\rho$, which then identifies $\delta$.
If Case 3(a) does not hold, $\frac{\boldsymbol{z}_{t+1}}{\boldsymbol{z}_{t}}+\frac{\boldsymbol{z}_{t}}{\boldsymbol{z}_{t+1}}$ must be constant in $t$, which is equivalent to the existence of some $\kappa\neq 0,-1,1$ such that
$r_{t}:=\frac{\boldsymbol{z}_{t+1}}{\boldsymbol{z}_{t}}\in\left\{\kappa,\frac{1}{\kappa}\right\}\ \text{for}\ t=1,2,3.$
This implies that if Case 3(a) does not hold, we can identify
$\frac{\sigma^{2}}{1+\rho}\left(1+\kappa+\frac{1}{\kappa}-\rho\right)$ (25)
from the equation above. Further using covariances, we identify
$\operatorname{Cov}\left(\frac{\Delta u_{i{t+2}}}{\boldsymbol{z}_{t+2}}-\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}},\frac{\Delta u_{i{t+1}}}{\boldsymbol{z}_{t+1}}-\frac{\Delta u_{i{t}}}{\boldsymbol{z}_{t}}\,|\,\delta_{i}=\delta\right)$
$=$ $\ \frac{1}{\boldsymbol{z}_{t+1}^{2}}\operatorname{Cov}\left(\frac{1}{r_{t+1}}\Delta u_{i{t+2}}-\Delta u_{i{t+1}},\Delta u_{i{t+1}}-r_{t}\Delta u_{i{t}}\,|\,\delta_{i}=\delta\right)$
$=$ $\ -\frac{1}{\boldsymbol{z}_{t+1}^{2}}\frac{\sigma^{2}}{1+\rho}\left(2+(1-\rho)\left(\frac{1}{r_{t+1}}+r_{t}-\frac{r_{t}}{r_{t+1}}\rho\right)\right).$ (26)
Now proceed to Cases 3(b) and 3(c).
Case 3(b): Assume that there are consecutive equal terms in $r_{t}$. Without loss of generality, assume that $r_{1}=r_{2}=\kappa$. Substituting $r_{1}$ and $r_{2}$ into (A.1) with $t=1$, the ratio of (A.1) to (25) identifies
$\frac{-\frac{1}{\boldsymbol{z}_{t+1}^{2}}\frac{\sigma^{2}}{1+\rho}\left(2+(1-\rho)\left(\frac{1}{r_{t+1}}+r_{t}-\frac{r_{t}}{r_{t+1}}\rho\right)\right)}{\frac{\sigma^{2}}{1+\rho}\left(1+\kappa+\frac{1}{\kappa}-\rho\right)}=C\frac{(1-\rho)(\kappa+1/\kappa-\rho)+2}{1+\kappa+1/\kappa-\rho},$ (27)
where $C\neq 0$ is a known constant. Define $\alpha:=\kappa+\frac{1}{\kappa}$, which satisfies $\alpha>2$ when $\kappa>0$, and $\alpha<-2$ when $\kappa<0$. From the expression above, we can identify
$\frac{(1-\rho)(\kappa+1/\kappa-\rho)+2}{1+\kappa+1/\kappa-\rho}=\rho+\frac{\alpha+2}{\rho-(\alpha+1)},$
whose value is denoted by $A$. If $\alpha<-2$, the function $\rho\mapsto\rho+{\alpha+2}/(\rho-(\alpha+1))$ is strictly increasing in $\rho\in(-1,1)$, which identifies $\rho$, and subsequently $\sigma^{2}$ from $\sigma^{2}/(1+\rho)(1+\kappa+1/\kappa-\rho)$. Otherwise, if $\alpha>2$, the equation
$\rho+\frac{\alpha+2}{\rho-(\alpha+1)}=A$ (28)
admits at most two distinct solutions $\rho_{1}$ and $\rho_{2}$, which are linked by
$\rho_{2}=\alpha+1+\frac{\alpha+2}{\rho_{1}-(\alpha+1)},$
and vice versa symmetrically. Suppose $\rho_{1}$ lies in (1,1)(-1,1). Consequently,
$\rho_{2}=\alpha+1+\frac{\alpha+2}{\rho_{1}-(\alpha+1)}\geq\alpha+1+\frac{\alpha+2}{1-(\alpha+1)}=\alpha-\frac{2}{\alpha}\geq 1,$
which necessarily falls outside (1,1)(-1,1). Hence, (28) yields at most one solution in (1,1)(-1,1), thereby identifying $\rho$, and subsequently $\delta$.
Case 3(c): Consider the case in which no consecutive terms in $(r_{1},r_{2},r_{3})$ are equal, i.e., $(r_{1},r_{2},r_{3})$ are alternating between $\kappa$ and $1/\kappa$ with $\kappa\neq\pm 1$. Without loss of generality, let $r_{1}=r_{3}=\kappa$ and $r_{2}=1/\kappa$. We observe that
$\boldsymbol{z}_{3}\boldsymbol{z}_{2}\operatorname{Cov}\left(\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}-\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}},\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}-\frac{\Delta u_{i{1}}}{\boldsymbol{z}_{1}}\,|\,\delta_{i}=\delta\right)$ $=-\frac{\sigma^{2}}{1+\rho}\left(\frac{2}{\kappa}+(1-\rho)\left(2-\kappa\rho\right)\right),$
$\boldsymbol{z}_{4}\boldsymbol{z}_{3}\operatorname{Cov}\left(\frac{\Delta u_{i{4}}}{\boldsymbol{z}_{4}}-\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}},\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}}-\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}\,|\,\delta_{i}=\delta\right)$ $=-\frac{\sigma^{2}}{1+\rho}\left(2\kappa+(1-\rho)\left(2-\rho/\kappa\right)\right).$
Taking the difference of these equations, we identify
$\left(\kappa-\frac{1}{\kappa}\right)\frac{\sigma^{2}}{1+\rho}\left(2+(1-\rho)\rho\right),$
where $\kappa-1/\kappa\neq 0$ is known. Additionally, we know the value of
$\boldsymbol{z}_{4}\boldsymbol{z}_{2}\operatorname{Cov}\left(\frac{\Delta u_{i{4}}}{\boldsymbol{z}_{4}}-\frac{\Delta u_{i{3}}}{\boldsymbol{z}_{3}},\frac{\Delta u_{i{2}}}{\boldsymbol{z}_{2}}-\frac{\Delta u_{i{1}}}{\boldsymbol{z}_{1}}\,|\,\delta_{i}=\delta\right)=\frac{\sigma^{2}}{1+\rho}(1-\rho)(\kappa\rho-1)(\rho-\kappa).$
Combined with (25) identified earlier, these two terms are used to identify the following:
$\frac{2+(1-\rho)\rho}{\rho-(\alpha+1)}=\rho+\alpha+\frac{(\alpha+2)(\alpha-1)}{\rho-(\alpha+1)}$
and
$\frac{(1-\rho)(\kappa\rho-1)(\rho-\kappa)}{\rho-(\alpha+1)}=\rho^{2}+\alpha-1+\frac{\alpha(\alpha+2)}{\rho-(\alpha+1)}.$
Since $\alpha$ is known, this information is equivalent to knowing $\rho+{(\alpha+2)(\alpha-1)}/({\rho-(\alpha+1)})$ and $\rho^{2}+{\alpha(\alpha+2)}/({\rho-(\alpha+1)})$, which can be used to identify
$\left[\rho^{2}+\frac{\alpha(\alpha+2)}{\rho-(\alpha+1)}\right]-\frac{\alpha}{\alpha-1}\left[\rho+\frac{(\alpha+2)(\alpha-1)}{\rho-(\alpha+1)}\right]=\rho^{2}-\frac{\alpha}{\alpha-1}\rho.$
Label the known values of $\rho+{(\alpha+2)(\alpha-1)}/({\rho-(\alpha+1)})$ and $\rho^{2}-\alpha\rho/(\alpha-1)$ as $B$ and $C$, respectively. Viewed as rational equations for $\rho$, this leads to the following system of quadratic equations:
$\rho^{2}-(B+1)\rho+B(\alpha+1)-2=0,$ (29)
$\rho^{2}-\frac{\alpha}{\alpha-1}\rho-C=0.$
We show by way of contradiction that (29) can be consistent with at most one $\rho$ in (1,1)(-1,1), thus uniquely determining $\rho$. For the quadratic system (29) to admit multiple solutions, note that the two equations must share the same roots, and hence coincide, i.e.,
$B=\frac{1}{\alpha-1},\quad C=1-\frac{2}{\alpha-1}.$
Plugging $C=1-\frac{2}{\alpha-1}$ into the second equation, the proof is complete once we verify that
$\rho^{2}-\frac{\alpha}{\alpha-1}\rho+\frac{3-\alpha}{\alpha-1}=0$
admits at most one solution in (1,1)(-1,1). To show this, evaluate this function at $\rho=1$:
$1-\frac{\alpha}{\alpha-1}+\frac{3-\alpha}{\alpha-1}=\frac{2-\alpha}{\alpha-1}.$
Similarly, at $\rho=-1$, we obtain
$1+\frac{\alpha}{\alpha-1}+\frac{3-\alpha}{\alpha-1}=\frac{\alpha+2}{\alpha-1}.$
Since $\alpha=\kappa+1/\kappa$ satisfies $|\alpha|>2$, we have
$\frac{2-\alpha}{\alpha-1}\frac{\alpha+2}{\alpha-1}=\frac{4-\alpha^{2}}{(\alpha-1)^{2}}<0,$
which shows that the function $\rho\mapsto\rho^{2}-\frac{\alpha}{\alpha-1}\rho+\frac{3-\alpha}{\alpha-1}$ takes values of opposite signs at $\rho=1$ and $\rho=-1$. Hence, it must have a unique root in (1,1)(-1,1), leading to a contradiction. This concludes Case 3(c).
Proof of Proposition 5.
We verify parts (i), (ii), and (iii) sequentially.
Proof of part (i): The regret consistency of $\hat{\sigma}^{2,{\mathrm{EB}}}_{i}$ and $\hat{\rho}^{{\mathrm{EB}}}_{i}$ is an immediate consequence of Theorem 3 because $(\sigma_{i}^{2},\rho_{i})$ have compact support by Assumption 1.
Proof of part (ii): We show that Assumption 4(ii) holds for $\tau_{i}\in\{a_{i},b_{i}\}$ with $p=2$. By Markov’s inequality, we have
$\int_{\Theta}\|\beta\|^{2}1\{\|\beta\|\geq M\}d\hat{G}_{N}(\theta)\leq\frac{1}{M^{\varepsilon}}\int_{\Theta}\|\beta\|^{2+\varepsilon}d\hat{G}_{N}(\theta).$
It follows from Lemma 5 that
$\int_{\Theta}\|\beta\|^{2}1\{\|\beta\|\geq M\}d\hat{G}_{N}(\theta)\leq\frac{1}{M^{\varepsilon}}\frac{A^{2+\varepsilon}}{N}\sum_{i=1}^{N}\|Y_{i}\|^{2+\varepsilon},$
and thus,
$\limsup_{N\to\infty}\mathbb{E}\left[\int_{\Theta}\|\beta\|^{2}1\{\|\beta\|\geq M\}d\hat{G}_{N}(\theta)\right]\leq\frac{A^{2+\varepsilon}}{M^{\varepsilon}}\mathbb{E}[\|Y_{i}\|^{2+\varepsilon}].$
One can make the right-hand side arbitrarily close to zero by choosing a sufficiently large $M$. Since $a_{i}$ and $b_{i}$ are dominated in absolute value by $\|\beta_{i}\|$, this proves Assumption 4(ii) for these parameters. It follows from Theorem 3 that $\hat{a}^{{\mathrm{EB}}}_{i}$ and $\hat{b}^{{\mathrm{EB}}}_{i}$ are regret-consistent.
Proof of part (iii): By Cauchy-Schwarz inequality, we observe that
$\frac{1}{N}\sum_{i=1}^{N}(\hat{Y}^{{\mathrm{EB}}}_{iT+1}-Y_{iT+1}^{*})^{2}/5$
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{2}|X_{2,iT+1}|^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{2}|Y_{iT}|^{2}+\frac{1}{N}\sum_{i=1}^{N}(\widehat{a\rho}_{i}^{{\mathrm{EB}}}-(a\rho)_{i}^{*})^{2}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}(\widehat{b\rho}_{i}^{{\mathrm{EB}}}-(b\rho)_{i}^{*})^{2}|X_{2,i,T}|^{2}$
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|Y_{iT}|^{4}}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}(\widehat{a\rho}_{i}^{{\mathrm{EB}}}-(a\rho)_{i}^{*})^{2}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\widehat{b\rho}_{i}^{{\mathrm{EB}}}-(b\rho)_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,i,T}|^{4}},$
where $(a\rho)_{i}:=a_{i}\rho_{i}$ and $(b\rho)_{i}:=b_{i}\rho_{i}$. Taking expectations on both sides, it suffices to show that the expectations of the terms on the RHS converge to zero.
Observe that $\tau_{i}\in\{a_{i},b_{i},a_{i}\rho_{i},b_{i}\rho_{i}\}$ are all dominated in absolute value by $\|\beta_{i}\|$ since $|\rho_{i}|\leq 1$. Combining the moment condition $\mathbb{E}[\|Y_{i}\|^{4+\varepsilon}]<\infty$ with the argument from Part (ii), we find that Assumption 4(ii) holds for $\tau_{i}\in\{a_{i},b_{i},a_{i}\rho_{i},b_{i}\rho_{i}\}$ with $p=4$. By Theorem 3 (with $p=2$), we have $\mathbb{E}[N^{-1}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}]\to 0$ and $\mathbb{E}[N^{-1}\sum_{i=1}^{N}(\widehat{a\rho}_{i}^{{\mathrm{EB}}}-(a\rho)_{i}^{*})^{2}]\to 0$ immediately. This addresses the first and fourth terms on the RHS. Applying Cauchy-Schwarz inequality, the second term on the RHS is bounded above by
$\mathbb{E}\left[\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}\right]\leq\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}\right]}\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}\right]}.$
Theorem 3 implies that $\mathbb{E}[N^{-1}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}]\to 0$, while $\mathbb{E}[N^{-1}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}]=\mathbb{E}[|X_{2,iT+1}|^{4}]$ is finite by assumption. Thus, the second term converges to zero. The third and fifth terms can be handled in an analogous way. We omit the details to avoid repetition. Each term on the right-hand side has expectation converging to zero. Hence $\mathbb{E}[N^{-1}\sum_{i=1}^{N}(\hat{Y}^{{\mathrm{EB}}}_{iT+1}-Y_{iT+1}^{*})^{2}]\to 0$ as desired, which establishes the regret consistency of $\hat{Y}_{iT+1}^{{\mathrm{EB}}}$.
A.2 Technical Lemmas
A.2.1 Identification Lemma (Theorem 1)
For the sake of completeness, we reproduce the key identification result from Bruni and Koch (1985) as Lemma 1, which is primarily used in this paper. Let $n\in\mathbb{N}$, $d\in\mathbb{N}$, and let $\Theta\subseteq\mathbb{R}^{n}$ denote a compact parameter space for $\theta$. Following the notation in Bruni and Koch (1985), we define the following class of functions:
$\Lambda(\Theta):=\left\{(\lambda_{1},\lambda_{2}):\Theta\to\mathbb{R}^{d}\times S_{+,d}\,{\Bigg|}\,\begin{aligned} &\ \|\lambda_{1}(\theta)\|+\|\accentset{\mbox{\large.}}{\lambda}_{1}(\theta)\|\leq K_{1},\\ &\ \|\lambda_{2}(\theta)\|+\|\accentset{\mbox{\large.}}{\lambda}_{2}(\theta)\|\leq K_{1},\ K_{2}{I}_{T}\leq\lambda_{2}(\theta)\leq K_{3}{I}_{T}\end{aligned}\right\},$
where $S_{+,d}$ denotes the set of $d\times d$ positive definite matrices, and $K_{j}$, $j=1,2,3$ are some positive constants. Each function $(\lambda_{1}(\theta),\lambda_{2}(\theta))$ represents the conditional mean and covariance of $Y_{i}\,|\,\theta_{i}\sim\mathcal{N}(\lambda_{1}(\theta_{i}),\lambda_{2}(\theta_{i}))$, where $\theta_{i}\in\Theta$ represents a random parameter. We also need the following restriction as in Bruni and Koch (1985): for any $(\lambda_{1},\lambda_{2}),(\tilde{\lambda}_{1},\tilde{\lambda}_{2})\in\Lambda(\Theta)$ and $\theta,\tilde{\theta}\in\Theta$, it holds that
$\lambda_{1}(\theta)=\tilde{\lambda}_{1}(\tilde{\theta}),\lambda_{2}(\theta)=\tilde{\lambda}_{2}(\tilde{\theta})\quad\Longrightarrow\quad\theta=\tilde{\theta}$. (30)
The condition in (30) requires that, for any $(\mu,\Sigma)\in\mathbb{R}^{d}\times S_{+,d}$, there exists at most one $\theta=\theta(\mu,\Sigma)\in\Theta$ such that
$\exists(\lambda_{1},\lambda_{2})\in\Lambda(\Theta):(\mu,\Sigma)=(\lambda_{1}(\theta),\lambda_{2}(\theta)),$
thus preventing multiple labels within $\Theta$. This condition is trivially satisfied when $\Lambda(\Theta)$ is a singleton consisting of a one-to-one function.
We write $\mathcal{M}(\Theta)$ for the set of complex Borel measures on $\Theta$, and $\mathcal{P}(\Theta)$ for its subset consisting of probability measures. Finally, define $\mathcal{T}:\Lambda(\Theta)\times\mathcal{M}(\Theta)\to C(\mathbb{R}^{d},\mathbb{C})$ by
$\mathcal{T}[\lambda_{1},\lambda_{2},\mu](y):=\int\frac{1}{(2\pi)^{d/2}|\lambda_{2}(\theta)|^{1/2}}\exp\left(-\frac{1}{2}\|\lambda_{2}(\theta)^{-1/2}(y-\lambda_{1}(\theta))\|^{2}\right)d\mu(\theta),$
where $C(\mathbb{R}^{d},\mathbb{C})$ denotes the class of all complex-valued continuous functions on $\mathbb{R}^{d}$. Note that, for $\mu\in\mathcal{P}(\Theta)$, $\mathcal{T}[\lambda_{1},\lambda_{2},\mu](x)$ represents the density of $\int_{\Theta}\mathcal{N}(\lambda_{1}(\theta),\lambda_{2}(\theta))d\mu(\theta)$, i.e., the marginal density of
$Y_{i}=\lambda_{1}(\theta_{i})+\lambda_{2}(\theta_{i})^{1/2}e_{i}$
under $\theta_{i}\sim\mu$ and $e_{i}\,|\,\theta_{i}\sim\mathcal{N}(0,I_{d})$. The following lemma from Bruni and Koch (1985) shows that $\mathcal{T}$ is an injective mapping, meaning that the mean and variance mixture components, as well as the mixing distribution, are identifiable from general Gaussian mixture distributions under the assumption of compact support.
Lemma 1 (Identification of Gaussian Mixture Models, Bruni and Koch, 1985)
Assume that $\mathcal{T}[\lambda_{1},\lambda_{2},\mu](y)=\mathcal{T}[\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\mu}](y)$ for all $y\in\mathbb{R}^{d}$, where $(\lambda_{1},\lambda_{2}),(\tilde{\lambda}_{1},\tilde{\lambda}_{2})\in\Lambda(\Theta)$ and $\mu,\tilde{\mu}\in\mathcal{P}(\Theta)$. Then, $\lambda_{1}=\tilde{\lambda}_{1}$, $\lambda_{2}=\tilde{\lambda}_{2}$, and $\mu=\tilde{\mu}$.
A.2.2 Consistency Lemmas (Theorem 2)
Let $\mathcal{P}_{\operatorname{sub}}$ denote the class of all subprobability measures on $\Theta$. We extend the metric $d$ to $\mathcal{P}_{\operatorname{sub}}$ by defining
$d(\mu_{0},\mu_{1}):=\sum_{r=1}^{\infty}\frac{1}{2^{r}}\left|\int_{\Theta}h_{r}(\theta)d\mu_{0}(\theta)-\int_{\Theta}h_{r}(\theta)d\mu_{1}(\theta)\right|,$ (31)
where $(h_{r})_{r=1}^{\infty}$ is the same as in Section 4. Let $\bar{\mathcal{G}}$ denote the closure of $\mathcal{G}$ relative to $\mathcal{P}_{\operatorname{sub}}$ with respect to the metric $d$. Since $d$ induces the vague topology on $\mathcal{P}_{\operatorname{sub}}$, and $\mathcal{P}_{\operatorname{sub}}$ is compact under this topology, $(\bar{\mathcal{G}},d)$ is itself a compact metric space (Kallenberg, 2021). We extend $f_{G}(Y,X)$, originally defined for $(G,Y,X)\in\mathcal{G}\times\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$, to $(\mu,Y,X)\in\bar{\mathcal{G}}\times\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$ by
$f_{\mu}(Y,X):=\int_{\Theta}\ell(Y\,|\,X,\theta)d\mu(\theta).$ (32)
The next lemma establishes that the extended likelihood is continuous in $\mu\in\bar{\mathcal{G}}$.
Lemma 2
Let Assumptions 1, 2, and 3 hold. Then, the mapping $\mu\mapsto f_{\mu}(Y,X):\bar{\mathcal{G}}\to\mathbb{R}$ is continuous for all $(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$.
Proof of Lemma 2.
Note that
$\ell(Y\,|\,X,\theta)=\frac{1}{(2\pi)^{T/2}|P(\delta)|}\exp\left(-\frac{1}{2}\|P(\delta)^{-1}(Y-X\beta)\|^{2}\right)\leq C_{1}\exp\left(-C_{2}\|Y-X\beta\|^{2}\right)$
for some constants $C_{1},C_{2}>0$ by Assumption 2(ii). Since $\operatorname{rank}(X)={d_{\beta}}$ for all $X\in\operatorname{supp}(X_{i})$ by Assumption 3(ii), it follows that $\|Y-X\beta\|^{2}\to\infty$ as $\|\beta\|\to\infty$, and hence $\lim_{\|\theta\|\to\infty}|\ell(Y\,|\,X,\theta)|=0$. Let $\varphi_{Y,X}(\theta):=\ell(Y\,|\,X,\theta)$. This implies that $\varphi_{Y,X}\in C_{0}(\Theta)$ for all pairs $(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$, where $C_{0}(\Theta)$ denotes the class of continuous functions $h:\Theta\to\mathbb{R}$ such that $\lim_{\|\theta\|\to\infty}|h(\theta)|=0$. It is well-known that $C_{0}(\Theta)$ is the completed space of $C_{c}(\Theta)$ with respect to the uniform norm $\|h\|_{\infty}=\sup_{\theta\in\Theta}|h(\theta)|$ (see, e.g., Section 21.4 of Royden and Fitzpatrick, 2010).
Now, consider a sequence $\mu_{n}\in\bar{\mathcal{G}}$ such that $\mu_{n}$ vaguely converges to $\mu\in\bar{\mathcal{G}}$, i.e., $\lim_{n\to\infty}d(\mu_{n},\mu)=0$. Fix a pair $(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$. Since $(h_{r})_{r=1}^{\infty}$ in (31) spans a dense subset of $C_{c}(\Theta)$, and hence of $C_{0}(\Theta)$, for any $\varepsilon>0$, there exists a real sequence $(a_{r})_{r=1}^{\bar{r}}$ with $\bar{r}<\infty$ such that $\left\|\varphi_{Y,X}-\sum_{r=1}^{\bar{r}}a_{r}h_{r}\right\|_{\infty}<\varepsilon$. This implies
$|f_{\mu_{n}}(Y,X)-f_{\mu}(Y,X)|$ $=\left|\int_{\Theta}\varphi_{Y,X}(\theta)d\mu_{n}(\theta)-\int_{\Theta}\varphi_{Y,X}(\theta)d\mu(\theta)\right|$
$\leq\left|\int_{\Theta}\sum_{r=1}^{\bar{r}}a_{r}h_{r}(\theta)d\mu_{n}(\theta)-\int_{\Theta}\sum_{r=1}^{\bar{r}}a_{r}h_{r}(\theta)d\mu(\theta)\right|+2\varepsilon$
$\leq\sup_{1\leq r\leq\bar{r}}|a_{r}2^{r}|\cdot d(\mu_{n},\mu)+2\varepsilon$
by construction of $d$, and hence
$\limsup_{n\to\infty}|f_{\mu_{n}}(Y,X)-f_{\mu}(Y,X)|\leq 2\varepsilon.$
Since $\varepsilon>0$ can be made arbitrarily small, it follows that $|f_{\mu_{n}}(Y,X)-f_{\mu}(Y,X)|\to 0$, completing the proof.
Lemma 3
Let Assumptions 1, 2, and 3 hold. Then, the following hold.
  1. (i)
    $\mathbb{E}_{G_{*}}[|\log f_{G_{*}}(Y_{i},X_{i})|]<\infty.$
  2. (ii)
    For any $\varepsilon>0$ and $\mu\in\bar{\mathcal{G}}$,
    $\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i}):=\sup\{f_{\tilde{\mu}}(Y_{i},X_{i}):d(\tilde{\mu},\mu)\leq\varepsilon\}$
    is a uniformly bounded random variable. Moreover,
    $\lim_{\varepsilon\to 0}\mathbb{E}[\log\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i})]=\mathbb{E}[\log f_{\mu}(Y_{i},X_{i})].$
  3. (iii)
    For any $\mu\in\bar{\mathcal{G}}$ with $\mu\neq G_{*}$, $f_{\mu}(Y_{i},X_{i})\neq f_{G_{*}}(Y_{i},X_{i})$ with positive probability.
Proof of Lemma 3.
Part (i): Since $\underaccent{\bar}{c}I_{T}\leq P(\delta)P(\delta)^{\prime}\leq\bar{c}I_{T}$ for all $\delta\in\mathcal{K}_{\delta}$ by Assumption 1(ii), we have
$C_{1}\int\exp(-C_{2}\|Y_{i}-X_{i}\beta\|^{2})dG_{*}(\theta)\leq f_{G_{*}}(Y_{i},X_{i})\leq C_{3}$
for some absolute constants $C_{1},C_{2},C_{3}>0$. Taking logarithms on both sides and applying Jensen’s inequality, it follows that
$\log(C_{1})-C_{2}\int\|Y_{i}-X_{i}\beta\|^{2}dG_{*}(\theta)\leq\log f_{G_{*}}(Y_{i},X_{i})\leq\log(C_{3}),$
which, in turn, implies
$|\log f_{G_{*}}(Y_{i},X_{i})|$ $\leq|\log(C_{1})|+|\log(C_{3})|+C_{2}\int\|Y_{i}-X_{i}\beta\|^{2}dG_{*}(\theta).$
Using $\|Y_{i}-X_{i}\beta\|^{2}\leq 2\|Y_{i}\|^{2}+2\|X_{i}\beta\|^{2}$, it follows that
$\mathbb{E}[|\log f_{G_{*}}(Y_{i},X_{i})|]\leq|\log(C_{1})|+|\log(C_{3})|+2\mathbb{E}[\|Y_{i}\|^{2}]+2\mathbb{E}_{G_{*}}[\|X_{i}\beta_{i}\|^{2}]<\infty,$
since $\mathbb{E}_{G_{*}}[\|X_{i}\beta_{i}\|^{2}]\leq\mathbb{E}_{G_{*}}[\|X_{i}\beta_{i}\|^{2}]+\mathbb{E}_{G_{*}}[\|P_{i}e_{i}\|^{2}]=\mathbb{E}[\|Y_{i}\|^{2}]<\infty$ by Assumption 3(i).
Part (ii): Let $\varepsilon>0$ and $\mu\in\bar{\mathcal{G}}$ be given. Since $(\bar{\mathcal{G}},d)$ is a compact metric space, $\{\tilde{\mu}\in\bar{\mathcal{G}}:d(\tilde{\mu},\mu)\leq\varepsilon\}$ is closed and hence compact, and therefore admits a countable dense subset $\{\tilde{\mu}_{j}\}_{j\geq 1}$. We thus have
$\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i})=\sup_{j\in\mathbb{N}}f_{\tilde{\mu}_{j}}(Y_{i},X_{i})$
by the continuity of $\mu\mapsto f_{\mu}$ established in Lemma 2. Since each $f_{\tilde{\mu}_{j}}(Y_{i},X_{i})$ is a random variable, we find that $\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i})$ is as well. The uniform boundedness of $\bar{f}_{\mu,\varepsilon}(Y_{i},X_{i})$ follows from that of $f_{\mu}(Y,X)$ across all $(Y,X)$ and $\mu\in\mathcal{P}_{\operatorname{sub}}$.
Part (iii): The case $\mu\in\mathcal{G}$ with $\mu\neq G_{*}$ is established in Theorem 1. Consider instead the case $\mu\in\bar{\mathcal{G}}\setminus\mathcal{G}$. Suppose, for contradiction, $f_{\mu}(Y_{i},X_{i})=f_{G_{*}}(Y_{i},X_{i})$ a.s. This implies
$\int_{\mathbb{R}^{T}}f_{\mu}(y,X_{i})dy=\int_{\mathbb{R}^{T}}\frac{f_{\mu}(y,X_{i})}{f_{G_{*}}(y,X_{i})}f_{G_{*}}(y,X_{i})dy=\mathbb{E}\left[\left.\frac{f_{\mu}(Y_{i},X_{i})}{f_{G_{*}}(Y_{i},X_{i})}\right|X_{i}\right]=1.$
This leads to a contradiction, since by Fubini’s theorem,
$\int_{\mathbb{R}^{T}}f_{\mu}(y,X_{i})dy=\int_{\Theta}\left(\int_{\mathbb{R}^{T}}\ell(y\,|\,X_{i},\theta)dy\right)d\mu(\theta)=\mu(\Theta)<1.$
A.2.3 Regret Consistency Lemmas (Theorem 3, Proposition 5)
Lemma 4
Let ${\tau}^{*}(Y,X)=\mathbb{E}_{G_{*}}[\tau_{i}\,|\,(Y_{i},X_{i})=(Y,X)]$ and $\hat{\tau}^{\mathrm{EB}}(Y,X)=\mathbb{E}_{\hat{G}}[\tau_{i}\,|\,(Y_{i},X_{i})=(Y,X)]$ denote the oracle and EB decision rules for $\tau_{i}$, respectively. Let $\mathcal{R}=\{(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i}):\|Y\|\leq M,\,M^{-1}{I}_{{d_{\beta}}}\leq X^{\prime}X\leq M{I}_{{d_{\beta}}}\}$ for a positive constant $M>1$. Let Assumptions 1, 2, 3, and 4 hold. Then, $\sup_{(Y,X)\in\mathcal{R}}|\hat{\tau}^{\mathrm{EB}}(Y,X)-{\tau}^{*}(Y,X)|\overset{\operatorname{a.s.}}{\longrightarrow}0$.
Proof of Lemma 4.
We first observe that
$\hat{\tau}^{\mathrm{EB}}(Y,X)=\frac{\int_{\Theta}\tau(\theta)\ell(Y\,|\,X,\theta)d\hat{G}(\theta)}{\int_{\Theta}\ell(Y\,|\,X,\theta)d\hat{G}(\theta)}=\frac{\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)d\hat{G}(\theta)}{\int_{\Theta}\varphi_{Y,X}(\theta)d\hat{G}(\theta)},$
where $\varphi_{Y,X}(\theta)=\ell(Y\,|\,X,\theta)$ is defined in the proof of Lemma 2. In that proof, we have established that $\varphi_{Y,X}\in C_{0}(\Theta)$ for each $(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$, and hence that
$\int_{\Theta}\varphi_{Y,X}(\theta)d\hat{G}(\theta)\overset{\operatorname{a.s.}}{\longrightarrow}\int_{\Theta}\varphi_{Y,X}(\theta)dG_{*}(\theta)$
as $d(\hat{G},G_{*})\overset{\operatorname{a.s.}}{\longrightarrow}0$. To further establish that the above convergence occurs uniformly in $(Y,X)\in\mathbb{R}^{T}\times\operatorname{supp}(X_{i})$, we verify that the family $\{\varphi_{Y,X}:(Y,X)\in\mathcal{R}\}$ forms a relatively compact subset of $C_{0}(\Theta)$ with respect to the sup-norm.
Since $X^{\prime}P(\delta)^{-1}P(\delta)^{-1}{}^{\prime}X\geq(M^{-1}\underaccent{\bar}{c}){I}_{T}$ for all $\delta\in\mathcal{K}_{\delta}$, by Assumption 1 and the construction of $\mathcal{R}$, there exist some positive constants $c_{1}$ and $c_{2}$ such that
$\sup_{(Y,X)\in\mathcal{R}}|\varphi_{Y,X}(\theta)|\leq c_{1}\exp\left(-c_{2}\|\beta\|^{2}\right).$
This implies that there exists $D>0$ such that
$\sup_{\theta:\|\beta\|\geq D}\sup_{(Y,X)\in\mathcal{R}}|\varphi_{Y,X}(\theta)|<\epsilon/2.$ (33)
Consider the restriction of $\varphi_{Y,X}$ to $B_{D}\times\mathcal{K}_{\delta}$, where $B_{D}:=\{\beta:\|\beta\|\leq D\}$. Arzela-Ascoli theorem implies that the class $\{\left.\varphi_{Y,X}\right|_{B_{D}\times\mathcal{K}_{\delta}}:(Y,X)\in\mathcal{R}\}$ is relatively compact in $C(B_{D}\times\mathcal{K}_{\delta})$ provided it is uniformly bounded and equi-continuous (Royden and Fitzpatrick, 2010). The uniform boundedness is obvious from the previous argument. The uniform equicontinuity follows from the fact that
$\sup_{(Y,X)\in\mathcal{R}}\left\|\nabla_{\theta}\varphi_{Y,X}(\theta)\right\|\leq c_{1}^{\prime}\exp\left(-c_{2}\|\beta\|^{2}\right)(\|\beta\|^{2}+1),$
for some $c_{1}^{\prime}>0$ and the same $c_{2}$ as before, where $\nabla_{\theta}=\partial/\partial\theta$. Taking the supremum over $\theta$, this implies that $\sup_{\theta\in B_{D}\times\mathcal{K}_{\delta}}\left(\sup_{(Y,X)\in\mathcal{R}}\left\|\nabla_{\theta}\varphi_{Y,X}(\theta)\right\|\right)<\infty$. Consequently, Arzela-Ascoli theorem implies that there exist $(Y_{1},X_{1}),\ldots,(Y_{K},X_{K})\in\mathcal{R}$, $K<\infty$ such that, for all $(Y,X)\in\mathcal{R}$,
$\min_{k=1,\ldots,K}\sup_{\theta\in B_{D}\times\mathcal{K}_{\delta}}|\varphi_{Y,X}(\theta)-\varphi_{Y_{k},X_{k}}(\theta)|<\epsilon$
from which, combined with (33), it follows
$\min_{k=1,\ldots,K}\sup_{\theta\in\Theta}|\varphi_{Y,X}(\theta)-\varphi_{Y_{k},X_{k}}(\theta)|<\epsilon.$
Therefore, $\{\varphi_{Y,X}\}_{(Y,X)\in\mathcal{R}}$ is relatively compact in $C_{0}(\Theta)$. The same line of arguments establish that $\{\tau\cdot\varphi_{Y,X}\}_{(Y,X)\in\mathcal{R}}$ is also relatively compact, whose details are omitted for brevity.
Since a relatively compact set admits a finite $\epsilon$-net for every $\epsilon>0$, by the standard approximation argument, we find that
$\sup_{(Y,X)\in\mathcal{R}}\left|\int_{\Theta}\varphi_{Y,X}(\theta)d\hat{G}(\theta)-\int_{\Theta}\varphi_{Y,X}(\theta)dG_{*}(\theta)\right|\overset{\operatorname{a.s.}}{\longrightarrow}0$
and
$\sup_{(Y,X)\in\mathcal{R}}\left|\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)d\hat{G}(\theta)-\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)dG_{*}(\theta)\right|\overset{\operatorname{a.s.}}{\longrightarrow}0.$
By the continuity of $f_{G_{*}}$ and the compactness of $\mathcal{R}$, we observe that
$\inf_{(Y,X)\in\mathcal{R}}\int_{\Theta}\varphi_{Y,X}(\theta)dG_{*}(\theta)=\inf_{(Y,X)\in\mathcal{R}}f_{G_{*}}(Y,X)\geq c>0$
for some $c>0$, and hence
$\liminf_{N\to\infty}\inf_{(Y,X)\in\mathcal{R}}\int_{\Theta}\varphi_{Y,X}(\theta)d\hat{G}(\theta)=\liminf_{N\to\infty}\inf_{(Y,X)\in\mathcal{R}}f_{\hat{G}}(Y,X)\geq c/2>0\quad\text{a.s.}$
Moreover, we have
$\sup_{(Y,X)\in\mathcal{R}}\left(\left|\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)dG_{*}(\theta)\right|+\left|f_{G_{*}}(Y,X)\right|\right)<\infty.$
This implies that
$\sup_{(Y,X)\in\mathcal{R}}\left|\hat{\tau}^{\mathrm{EB}}(Y,X)-{\tau}^{*}(Y,X)\right|$ $=\sup_{(Y,X)\in\mathcal{R}}\left|\frac{\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)d\hat{G}(\theta)}{f_{\hat{G}}(Y,X)}-\frac{\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)dG_{*}(\theta)}{f_{G_{*}}(Y,X)}\right|$
$\leq\sup_{(Y,X)\in\mathcal{R}}\left|\frac{\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)d\hat{G}(\theta)-\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)dG_{*}(\theta)}{f_{\hat{G}}(Y,X)}\right|$
$\quad+\sup_{(Y,X)\in\mathcal{R}}\left|\frac{\int_{\Theta}\tau(\theta)\varphi_{Y,X}(\theta)dG_{*}(\theta)}{f_{G_{*}}(Y,X)}\frac{f_{\hat{G}}(Y,X)-f_{G_{*}}(Y,X)}{f_{\hat{G}}(Y,X)}\right|$
$\overset{\operatorname{a.s.}}{\longrightarrow}0.$
Lemma 5
Assume the HIVDX model in (1). Let Assumption 1 hold. Further suppose that $|\bar{X}_{2,i}|\leq M$ and $\sum_{t=1}^{T}(X_{2,it}-\bar{X}_{2,i})^{2}\geq c$ for some constants $M,c>0$ and all $i=1,\ldots,N$. Then, there exists an absolute constant $A<\infty$ depending only on $\bar{c}$, $\underaccent{\bar}{c}$, $M$, $c$, and $T$, such that
$\left(\int_{\Theta}\|\beta\|^{p}d\hat{G}_{N}(\theta)\right)^{1/p}\leq A\left(\frac{1}{N}\sum_{i=1}^{N}\|Y_{i}\|^{p}\right)^{1/p}\quad\text{for all }p\geq 2,$
where $\hat{G}_{N}$ denotes the NPMLE obtained from the sample $(Y_{i},X_{i})_{i=1}^{N}$.
Proof of Lemma 5.
By the first-order conditions for the NPMLE, we have
$\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\leq 1\quad\text{for all }\theta\in\Theta,$
and $\hat{G}\large(\{\theta\in\Theta:N^{-1}\sum_{i=1}^{N}{\ell(Y_{i}\,|\,X_{i},\theta)}/f_{\hat{G}}(Y_{i},X_{i})=1\}\large)=1$. This implies
$\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\frac{\partial\log\ell(Y_{i}\,|\,X_{i},\theta)}{\partial\beta}=0$
for $\hat{G}-$a.e. $\theta$, and hence
$\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}X_{i}^{\prime}(P(\delta)P(\delta)^{\prime})^{-1}(Y_{i}-X_{i}\beta)=0.$
Rearranging the equation above, we obtain
$\beta=\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}X_{i}^{\prime}(P(\delta)P(\delta)^{\prime})^{-1}X_{i}\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}X_{i}^{\prime}(P(\delta)P(\delta)^{\prime})^{-1}Y_{i}\right).$
for $\hat{G}-$a.e. $\theta$. Regarded as a weighted regression of $P(\delta)^{-1}Y_{i}$ on $P(\delta)^{-1}X_{i}$, this yields the following inequality:
$\beta^{\prime}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}X_{i}^{\prime}(P(\delta)P(\delta)^{\prime})^{-1}X_{i}\right)\beta\leq\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\|P(\delta)^{-1}Y_{i}\|^{2}.$
By Assumption 1, we have $\bar{c}^{-1}{I}_{T}\leq(P(\delta)P(\delta)^{\prime})^{-1}\leq\underaccent{\bar}{c}^{-1}{I}_{T}$. Let $\mu_{i}=(1/T)\sum_{t=1}^{T}X_{2,it}$ and $s_{i}^{2}=(1/T)\sum_{t=1}^{T}(X_{2,it}-\mu_{i})^{2}$. Then, the smallest eigenvalue of $X_{i}^{\prime}X_{i}$ satisfies
$\lambda_{\min}(X_{i}^{\prime}X_{i})\geq\frac{Ts_{i}^{2}}{1+\mu_{i}^{2}+s_{i}^{2}}\geq\underaccent{\bar}{\lambda}>0\quad\text{for all }i=1,,\ldots,N,$
where $\underaccent{\bar}{\lambda}=\underaccent{\bar}{\lambda}(M,c)={c}/({1+M^{2}+c/T})$ is a positive constant. This implies that
$\|\beta\|^{2}=\|\beta\|^{2}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\leq\frac{\bar{c}}{\underaccent{\bar}{c}\underaccent{\bar}{\lambda}}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\|Y_{i}\|^{2}=A^{2}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\|Y_{i}\|^{2}$
where $A^{2}={\bar{c}}/(\underaccent{\bar}{c}\underaccent{\bar}{\lambda}(M,c))$. By the Jensen’s inequality, it follows that, for $p\geq 2$,
$\|\beta\|^{p}\leq A^{p}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\|Y_{i}\|^{2}\right)^{p/2}\leq A^{p}\frac{1}{N}\sum_{i=1}^{N}\frac{\ell(Y_{i}\,|\,X_{i},\theta)}{f_{\hat{G}}(Y_{i},X_{i})}\|Y_{i}\|^{p}.$
Integrating both sides with respect to $\hat{G}$ yields
$\int_{\Theta}\|\beta\|^{p}d\hat{G}(\theta)\leq A^{p}\frac{1}{N}\sum_{i=1}^{N}\|Y_{i}\|^{p}.$
Appendix B Auxiliary Identification Results
B.1 Verification of Assumption 2 for ARMA(1,1) Errors
We specify the choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in the context of the HIVD model with ARMA(1,1) errors. For simplicity, we assume $T=4$. To illustrate the construction, let $\boldsymbol{X}$} be a column of ones (corresponding to the random intercept) and define the differencing matrix
$\boldsymbol{M}=\begin{pmatrix}-1&0&0\\ 1&-1&0\\ 0&1&-1\\ 0&0&1\end{pmatrix}$.
This yields the differenced vector $\boldsymbol{M}^{\prime}Y_{i}=(\Delta u_{i2},\Delta u_{i3},\Delta u_{i4})^{\prime}$. Consequently, the transformed covariance matrix $\mathcal{V}:=\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}=\mathbb{V}(\boldsymbol{M}^{\prime}Y_{i}\mid\delta_{i}=\delta)$ has elements:
$\mathcal{V}_{11}$ $=2\,\frac{1+\varphi^{2}-\varphi(1-\rho)}{1+\rho}\,\sigma^{2},$
$\mathcal{V}_{12}$ $=\frac{\varphi^{2}(\rho-1)+\varphi(\rho^{2}-\rho+2)+(\rho-1)}{1+\rho}\,\sigma^{2},$
$\mathcal{V}_{13}$ $=\frac{(\rho-1)(\rho+\varphi)(1+\rho\varphi)}{1+\rho}\,\sigma^{2}.$
Using these moments, $\rho$ is explicitly identified as
$\rho=1+\frac{\mathcal{V}_{13}}{\mathcal{V}_{11}/2+\mathcal{V}_{12}},$
where the condition $\rho+\varphi\neq 0$ guarantees that the denominator is non-zero. Once $\rho$ is determined, $\varphi$ and $\sigma^{2}$ are recovered from the following system:
$\frac{1+\rho}{2}\mathcal{V}_{11}+(1-\rho)(\rho\mathcal{V}_{12}-\mathcal{V}_{13})$ $=(1+\varphi^{2})\sigma^{2},$ (34)
$\rho\mathcal{V}_{12}-\mathcal{V}_{13}$ $=\varphi\sigma^{2}.$ (35)
We distinguish two cases based on the left-hand side of (35):
  1. 1.
    If $\rho\mathcal{V}_{12}-\mathcal{V}_{13}=0$, then $\varphi=0$ (since $\sigma^{2}>0$). In this scenario, $\sigma^{2}$ is directly identified by (34).
  2. 2.
    If $\rho\mathcal{V}_{12}-\mathcal{V}_{13}\neq 0$, then $\varphi\neq 0$. We take the ratio of (34) to (35) to eliminate $\sigma^{2}$:
    $\frac{1+\varphi^{2}}{\varphi}=\frac{\frac{1+\rho}{2}\mathcal{V}_{11}+(1-\rho)(\rho\mathcal{V}_{12}-\mathcal{V}_{13})}{\rho\mathcal{V}_{12}-\mathcal{V}_{13}}\equiv\mathcal{R}.$
    Rearranging this expression leads to the quadratic equation $\varphi^{2}-\mathcal{R}\varphi+1=0$. From the properties of quadratic equations, $\mathcal{R}$ represents the sum of the two roots, while the constant $1$ represents their product. This implies that the roots are reciprocals, say $z$ and $1/z$. The condition $|\mathcal{R}|\geq 2$ (which holds by construction for any real $\varphi$) ensures that the roots are real. Finally, the assumption $|\varphi|\leq 1$ allows us to uniquely identify $\varphi$ as the root that is less than or equal to 1 in absolute value.
Having identified $\varphi$, $\sigma^{2}$ is determined by $\sigma^{2}=(\rho\mathcal{V}_{12}-\mathcal{V}_{13})/\varphi$. Therefore, we have verified the required one-to-one mapping between $\mathcal{V}$ and $\delta$.
B.2 Identification Failure of the HIVDX model with $T=4$
Proposition 3 indicates that $G_{*}$ is not identified when $T=4$ and the sequence of first differences $(\Delta X_{2,it})_{t=2}^{4}$ takes the symmetric form
$(\Delta X_{2,i2},\Delta X_{2,i3},\Delta X_{2,i4})=(x,\tilde{x},x),$
with $|\tilde{x}|\neq|x|$. To see this, consider the case where $X_{2,it}=\mathbf{1}\{t\geq 3\}$ for all $i$:
$Y_{it}=a_{i}+b_{i}\mathbf{1}\{t\geq 3\}+u_{it},\quad t=1,\ldots,4,$ (36)
with $u_{it}$ following the structure in (1b). Here, the difference sequence is $(0,1,0)$. To eliminate the random slope $b_{i}$ (which appears only in the difference $\Delta Y_{i3}$), we construct $\boldsymbol{M}$ to select only the second and fourth differences:
$\boldsymbol{M}=\begin{pmatrix}-1&0\\ 1&0\\ 0&-1\\ 0&1\end{pmatrix}$.
This yields $\boldsymbol{M}^{\prime}Y_{i}=(\Delta u_{i2},\Delta u_{i4})^{\prime}=(u_{i2}-u_{i1},u_{i4}-u_{i3})^{\prime}$. The corresponding covariance matrix $\mathcal{V}:=\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}=\mathbb{V}(\boldsymbol{M}^{\prime}Y_{i}\mid\delta_{i}=\delta)$ is $2\times 2$ and imposes the following two restrictions on $\delta=(\sigma^{2},\rho)$:
$\mathcal{V}_{11}$ $=\mathcal{V}_{22}=\frac{2\sigma^{2}}{1+\rho},$
$\mathcal{V}_{12}$ $=\mathcal{V}_{21}=-\frac{\rho(1-\rho)\sigma^{2}}{1+\rho}.$
These relations are insufficient to uniquely determine $\delta$. For instance, the distinct parameter pairs $(1.04,0.2)$ and $(1.56,0.8)$ yield identical covariance matrices $\mathcal{V}$.
The intuition for this identification failure is straightforward. Because the differenced regressor sequence $(0,1,0)$ is symmetric and the error process $u_{it}$ is stationary, the joint distribution of the utilized differences $(\Delta Y_{i2},\Delta Y_{i4})$ is symmetric. Specifically, $\mathcal{V}_{11}=\mathcal{V}_{22}$ provides the scale of the variance, while the correlation depends on the ratio $\mathcal{V}_{12}/\mathcal{V}_{11}=-\rho(1-\rho)/2$. Since the function $f(\rho)=\rho(1-\rho)$ is not injective on (1,1)(-1,1), solving for $\rho$ involves a quadratic equation with potentially two distinct valid roots, precluding point identification.
This failure highlights that sufficient time variation in the covariates is a prerequisite for identification in short panels. However, as established in Proposition 4, the additional variation afforded by a fifth time period breaks this symmetry and ensures that $\delta$ is uniquely determined.
Figure 4: Pairwise scatter plots of $\hat{\theta}_{i}^{\mathrm{EB}}$
(a) Pair $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})$.
(b) Pair $(\hat{a}_{i}^{\mathrm{EB}},\hat{\sigma}_{i}^{2,\mathrm{EB}})$.
(c) Pair $(\hat{b}_{i}^{\mathrm{EB}},\hat{\sigma}_{i}^{2,\mathrm{EB}})$.
(d) Pair $(\hat{a}_{i}^{\mathrm{EB}},\hat{\rho}_{i}^{\mathrm{EB}})$.
(e) Pair $(\hat{b}_{i}^{\mathrm{EB}},\hat{\rho}_{i}^{\mathrm{EB}})$.
(f) Pair $(\hat{\rho}_{i}^{\mathrm{EB}},\hat{\sigma}_{i}^{2,\mathrm{EB}})$.
Note: Each panel shows the scatter plot of a pair of components of $(\hat{\theta}_{i}^{\mathrm{EB}})_{i=1}^{N}$.
Appendix C Additional Empirical Results
Figure 4 shows the pairwise scatter plots of the EB estimates for $\theta_{i}$ computed from the sample of $N=938$ individuals, supporting the heterogeneity pattern reported in Table 1.151515To be precise, by the law of total variance, $\mathbb{V}(\theta_{i})=\mathbb{V}(\hat{\theta}_{i}^{\mathrm{EB}})+\mathbb{E}[\mathbb{V}(\theta_{i}|(Y_{it},X_{2,it})_{t=1}^{T})]$, where the second term represents the expected posterior variance, reflecting the average variance of the EB estimator. Panel (a) illustrates that individual intercepts ($\hat{a}_{i}^{\mathrm{EB}}$) and slopes ($\hat{b}_{i}^{\mathrm{EB}}$) are negatively correlated. Panels (b) and (c) also reveal negative associations between $\hat{a}_{i}^{\mathrm{EB}}$ and $\hat{\sigma}_{i}^{2,\mathrm{EB}}$, and between $\hat{b}_{i}^{\mathrm{EB}}$ and $\hat{\sigma}_{i}^{2,\mathrm{EB}}$, respectively. In the remaining panels, individual persistence levels appear only weakly correlated with the other individual-specific parameters.